### v1

# MultiBanana: A Challenging Benchmark for Multi-Reference Text-to-Image Generation

Abstract
Recent text-to-image generation models have acquired the ability of multi-reference generation and editing; the ability to inherit the appearance of subjects from multiple reference images and re-render them under new contexts. However, the existing benchmark datasets often focus on the generation with single or a few reference images, which prevents us from measuring the progress on how model performance advances or pointing out their weaknesses, under different multi-reference conditions. In addition, their task definitions are still vague, typically limited to axes such as “what to edit” or “how many references are given”, and therefore fail to capture the intrinsic difficulty of multi-reference settings. To address this gap, we introduce MultiBanana, which is carefully designed to assesses the edge of model capabilities by widely covering multi-reference-specific problems at scale: (1) varying the number of references, (2) domain mismatch among references (e.g., photo vs. anime), (3) scale mismatch between reference and target scenes, (4) references containing rare concepts (e.g., a red banana), and (5) multilingual textual references for rendering. Our analysis among a variety of text-to-image models reveals their superior performances, typical failure modes, and areas for improvement. MultiBanana will be released as an open benchmark to push the boundaries and establish a standardized basis for fair comparison in multi-reference image generation. Our data and code are available at https://github.com/matsuolab/multibanana.
Figure 1: The overview of MultiBanana. MultiBanana widely covers multi-reference specific problems, varying the number of references (the top row), domain and scale mismatch among references (two on the left in the middle row), multilingual text rendering (center in the bottom row), and containing rare concepts (right in the bottom row).
[…]
1 Introduction
Recent advances in image generation, driven by large-scale multimodal LLM backbones, have enabled unprecedented levels of instruction-following [google2025nanobanana, openai2025gpt4oimage, wu2025qwenimagetechnicalreport, cao2025hunyuanimage3, wu2025less, xiao2024omnigen, wu2025omnigen2, xia2025dreamomni, xia2025dreamomni2, cui2025emu35nativemultimodalmodels]. Among these developments, multi-reference image generation has emerged as a particularly notable capability. When a user provides multiple reference images, the model can inherit the subject’s appearance and re-render it in a new context while faithfully preserving identity. This shift moves image generation beyond single-reference conditioning toward greater controllability, driving increasing demand in industrial domains such as content production [ruiz2022dreambooth, ye2023ip-adapter], advertising [horita2024retrievalaugmented, morita2025tkg], and fashion design [zhu2023tryondiffusion, choi2024improving, fang2024vivid], where reference-guided generation is essential.
However, despite this progress, we lack a benchmark that systematically evaluates how model performance varies under different multi-reference conditions. As shown in Table 1, existing benchmarks [wang2023editbench, basu2023editval, sheynin2024emuedit, zhang2023magicbrush, huang2024smartedit, yu2025anyedit, ma2024i2ebench, ye2025imgedit, ruiz2022dreambooth, wu2025omnigen2, xia2025dreamomni2] suffer from two fundamental limitations. First, they have an extremely narrow coverage due to the small number of unique reference images, making it impossible to assess the impact of varying reference conditions in both single- and multi-reference tasks. Second, their task definitions still lack details, typically limited to axes such as“what to edit” or “how many references are given”, and therefore fail to capture the intrinsic difficulty of multi-reference settings. Moreover, existing benchmarks limit the number of references to a narrow range (typically 1–5) and do not examine how performance changes with more references.
To address these limitations, we introduce a new evaluation framework grounded not only in task categories but also in the intrinsic properties of the reference set itself. Our benchmark incorporates multiple orthogonal axes specific to the multi-reference regime, including broader coverage of the number of references (up to 8), domain gaps between references (e.g., photo–anime mixtures), mismatches in object scale between reference and target scenes, rare concepts in the references (e.g., a red banana), and multilingual textual rendering references (English, Chinese, and Japanese).
Using this benchmark, we conduct a cross-conditional analysis to assess model capabilities, identify characteristic failure modes, and point out remaining gaps. For example, while performance degradation due to increasing reference counts can be mitigated by agent-style multi-stage generation pipelines, failures arising from unsupported styles or languages remain difficult to remedy with inference-time strategies. We will open-source MultiBanana to establish a standard foundation for fair comparison and to accelerate progress in multi-reference image generation.
Table 1: Comparison among major benchmarks for reference-based image generation. Existing benchmarks do not provide systematic evaluation across diverse multi-reference conditions, support only a limited number of references, and fail to adequately account for important factors such as differences and compatibility among reference images. Our benchmark expands the upper limit on the number of references and introduces difficult reference combinations, including domain mismatches, scale mismatches, rare concepts, and multilingual prompts, thereby explicitly addressing challenges unique to multi-reference image generation.
Benchmark #Size #Tasks #References Difficult Reference Combination Metrics
EditBench [wang2023editbench] 240 1 1 CLIP [radford2021clip]
EditVal [basu2023editval] 648 13 1 CLIP, VLM, manual
EmuEdit [sheynin2024emuedit] 3055 7 1 L1, CLIP, DINO [caron2021dino]
MagicBrush [zhang2023magicbrush] 1053 9 1 L1, L2, CLIP, DINO
AnyEdit [yu2025anyedit] 1250 25 1 L1, CLIP, DINO
I2EBench [ma2024i2ebench] 2240 16 1 GPT [openai2023gpt4]
ImgEdit-Bench [ye2025imgedit] 811 14 1 GPT (3 dim.), Fake Det. [xu2024fakeshield]
DreamBooth [ruiz2022dreambooth] 75 1 1 CLIP. DINO
OmniContext [wu2025omnigen2] 400 8 3 GPT (3 dim.)
DreamOmni2 [xia2025dreamomni2] 319 29 4 Gemini [geminiteam2023gemini], Doubao [bytedance2025doubao]
MultiBanana (Ours) 3769 36 8 GPT, Gemini (5 dim.)
2 Related Works
2.1 Controllable Text-to-Image Generation
Diffusion models have achieved state-of-the-art performance in high-fidelity image synthesis. Recent large pretrained diffusion models with textual conditioning, such as Stable Diffusion [rombach2022ldm, podell2024sdxl, esser2024sd3] and FLUX [flux2024, labs2025flux1kontext], enable the synthesis of novel scene images while preserving the identity depicted in the reference images. Building on these developments, approaches based on fine-tuning (e.g., DreamBooth [ruiz2022dreambooth]) and adapter-based methods (e.g., IP-Adapter [ye2023ip-adapter]) have been actively explored to address the controllability [jiang2025infiniteyou, gal2023an, brooks2023instructpix2pix, mokady2023null, tan2025ominicontrol, NEURIPS2023_602e1a5d]. Additionally, training-free approaches (e.g., prompt-to-prompt [hertz2023prompttoprompt]) have also been widely proposed [parmar2023zero, Cao_2023_ICCV, Huberman-Spiegelglas_2024_CVPR, Wallace_2023_CVPR, Tumanyan_2023_CVPR, kulikov2025flowedit, miyake2025negative-prompt-inversion]. Recent research on unified image generation has gained momentum, aiming to handle multiple image generation tasks within a single model. For example, ChatGPT-Image-1 [openai2025gpt4oimage] and Nano Banana [google2025nanobanana] can jointly process text and image inputs, enabling integrated image generation and editing within a unified framework. Similarly, the open-source models, such as Qwen-Image [wu2025qwenimagetechnicalreport] and FLUX.1-Kontext-dev [labs2025flux1kontext], built upon a robust multimodal large language model backbone, demonstrates high-quality and flexible image generation and editing performance. Beyond these, numerous open-source unified generative models continue to emerge [xiao2024omnigen, xia2025dreamomni, wu2025omnigen2, xia2025dreamomni2]. These advancements collectively demonstrate that diffusion models are becoming increasingly flexible and adaptable across diverse conditional generation settings.
2.2 Benchmarks for Reference-Based Generation
Among reference-driven image generation tasks, image editing can be the most widely recognized, which has referred to the task of generating an edited image based on a single reference image and a textual instruction describing how the edit should be performed. Representative benchmarks such as MagicBrush [zhang2023magicbrush] and EMU-Edit [sheynin2024emuedit] have extended task-specific evaluations but still rely heavily on similarity-based metrics. SmartEdit [huang2024smartedit] supports editing of complex scenes, but does not sufficiently cover more general settings. Similarly, I2E-Bench [ma2024i2ebench] employs GPT-4o to provide human-aligned evaluations across a variety of editing tasks, yet it uses different metrics for each task and therefore fails to capture the shared characteristics of editing as a whole. More recently, ImgEdit [ye2025imgedit] has enabled the evaluation of multi-turn editing tasks, where image generation and editing are alternated interactively. One of the earliest benchmarks for reference-based image generation beyond instruction-based editing is the DreamBooth [ruiz2022dreambooth] dataset. However, such benchmarks are still limited to generation tasks conditioned on a single reference image and thus cannot handle more complex or compositional settings. Recently, following advances demonstrated by GPT-Image-1 [openai2025gpt4oimage] and Nano Banana [google2025nanobanana], increasing attention has been directed toward multi-reference image generation, where multiple reference images are used jointly for generation. Benchmarks such as OmniContext [wu2025omnigen2] and DreamOmni2 [xia2025dreamomni2] have been proposed to evaluate this task. Nevertheless, these benchmarks encompass only a small number of images and task types, thereby limiting the scope of the evaluation. Moreover, despite being designed for multi-reference generation, they do not evaluate qualitative differences or relationships between reference images, and their setups remain similar to those of conventional single-image editing benchmarks.
Figure 2: Construction pipeline for our benchmark, consisting of four stages: (1) collecting high-quality real and synthetic images, (2) filtering out inappropriate or low-quality samples, (3) performing hierarchical category classification, and (4) generating and validating editing instructions by Gemini and humans.
3 MultiBanana
3.1 Task Types
In reference-driven image generation, tasks can be categorized based on the number and relational structure of reference images. In this work, we follow the overall design philosophy of prior benchmarks such as DreamOmni2 [xia2025dreamomni2], but place a stronger emphasis on an aspect that has been largely underexplored so far: the diversity and compositional relationships among multiple references. While existing benchmarks mainly evaluate performance in terms of the number of references or single-image editing accuracy, our benchmark explicitly aims to measure a model’s ability to understand and reason about inter-reference variations as a key part of multimodal generation.
Single Reference.   This is the most fundamental setting, corresponding to the standard image editing task. Given a single reference image and a textual instruction, the model generates an edited image that maintains semantic consistency and preserves the subject’s visual identity. This configuration follows traditional instruction-driven editing benchmarks such as MagicBrush [zhang2023magicbrush] and EMU-Edit [sheynin2024emuedit], which represent the foundation of reference-driven generation.
Two References – 11 Task Types.   The two-reference setting provides the model with two reference images (typically a main subject and an auxiliary one) and a textual instruction. We define 11 editing task types, following prior work, while balancing feasibility and challenge: subject addition, subject replacement, background change, color modification, material modification, pose modification, hairstyle modification, makeup modification, tone transformation (lighting, season, weather, etc.), style transfer (oil painting, sketch, anime, etc.), and text correction (in-image text editing or replacement). These tasks span both local attribute manipulations and global transformations, evaluating a model’s ability to maintain visual coherence, subject identity, and contextual adaptability.
Multi References – Compositional Generation and Inter-Reference Relationships.   The multi-reference setting uses 3–8 reference images, requiring the model to perform more compositional and reasoning-based generation. While prior works, such as DreamOmni2 [xia2025dreamomni2], mainly focus on compositional generation using multiple references, our benchmark takes a step further by making inter-reference diversity, consistency, and structural relations central to the evaluation.
We define four base task structures, each evaluated with $\{3,4,5,6,7,8\}$ reference images, and include variants with and without textual instructions — resulting in a total of 48 tasks:
  • X Objects: Compose $X$ objects from the reference set.
  • X–1 Objects + Background: Place X1X-1 objects within the background of another reference.
  • X–1 Objects + Local: Apply a local change (specified by a reference) to one of the X1X-1 objects.
  • X–1 Objects + Global: Arrange X1X-1 objects and modify the overall tone according to another reference.
This design enables us to assess whether a model can effectively integrate multiple reference sources into a coherent, contextually consistent scene.
Diversity Factors in Multi-Reference Generation.   A central contribution of our benchmark is the explicit incorporation of inter-reference diversity factors, which have been largely overlooked in previous benchmarks such as DreamOmni2. We highlight the following new evaluation dimensions:
  • Number of references: systematic comparison across 3–8 inputs.
  • Cross-domain diversity: mixing references from different domains such as real photos, anime, CG, or sketches.
  • Scale and viewpoint differences: variation in object size or camera perspective between references and outputs.
  • Rare concept references: inclusion of uncommon concepts (e.g., “a red banana”).
  • Multilingual references: assessing multilingual instruction understanding during rendering.
These factors go beyond conventional image-editing evaluation and are critical for assessing a model’s compositional reasoning, identity preservation, and cross-domain generalization. In summary, our benchmark diverges from the conventional notion of multi-reference generation, such as in DreamOmni2, which primarily focuses on simple reference composition. Instead, it introduces a new framework designed to evaluate model capabilities to interpret relationships among heterogeneous references and integrate them into coherent generations.
Figure 3: (Left) Comparison between the statistics of real data only and those after adding synthetic data. The original dataset was biased toward background images, with few person- and object-related samples. To correct this imbalance, we generated additional synthetic images using Nanobanana and ChatGPT-Image-1, focusing on clear subjects such as people, animals, and objects. This significantly increased person- and object-related categories, resulting in a more balanced and comprehensive benchmark. (Right) Examples of synthesized images in each category.
3.2 How to Construct MultiBanana
Image Collection.   The reference images used in our benchmark consist of both real and synthetic data. All real images were extracted from the LAION-5B dataset [schuhmann2022laion]. To ensure high quality, we selected only images with an Aesthetic score higher than 6.25 and a resolution greater than 512 pixels. However, the collected real images exhibited a distributional bias (Figure 3; left). Specifically, landscape scenes appeared frequently, whereas categories useful as editing references, such as humans and objects, were underrepresented. While landscape images can serve as useful background references, the scarcity of human and object references limits the diversity of editing tasks that can be constructed in the benchmark. To address this imbalance, we generated additional synthetic images under diverse conditions using Nano Banana [google2025nanobanana] and ChatGPT-Image-1 [openai2025gpt4oimage], and used them as supplementary reference images (Figure 3; right). This approach not only mitigated the categorical bias in the reference set but also expanded the coverage and diversity of reference images by incorporating both real and synthetic examples.
Image Filtering.  Among the collected images, some contained small subjects that were difficult to use as references, or were inappropriate as reference images, such as those containing harmful content, tables, or corrupted data. To filter out such images and retain only those suitable for a reference-based image editing benchmark, we applied the following procedure. Following [ye2025imgedit], we extracted images that contained large and clearly visible objects suitable for editing tasks. Object detection was performed using YOLOv12 [tian2025yolov12] and SAM2 [ravi2025sam], while CLIP [radford2021clip] was used to verify semantic consistency. Additionally, harmful or corrupted images were removed through a combination of automatic filtering with Gemini and manual inspection, resulting in the final curated set of real images.
Figure 4: (Left) Breakdown of the multi-reference tasks. The editing sets were selected to ensure that the number of sets within each task is balanced across different reference counts. (Middle) For every X–references task, the dataset contains at least 390 editing sets. Further, each colored task category also includes at least 70 sets, which is larger than the prior work [xia2025dreamomni2]. (Right) Word cloud generated from all prompts. It primarily consists of terms that describe a wide range of object categories as well as words indicating spatial directions.
Category Classification. To construct multi-reference image generation tasks, we categorized each reference image accurately. For example, in a portrait transformation task that modifies a person’s hairstyle or makeup, both reference images must contain a person. Similarly, in replacement tasks involving humans or objects, both reference images must include either a person or an object, and for background replacement tasks, images categorized as background must be used. To satisfy such task-dependent conditions, we performed hierarchical image classification on a combined set of real and synthetic images. We first obtained automatic multi-level labels using Gemini and then conducted human verification to ensure correct categorization. At the top level, we defined five major categories: person, object, background, light, tone. Each of these top-level categories was further divided into more specific subcategories. For instance, under the person category, we included male, female, hairstyle, makeup, pose useful, and face expression; and under the object category, we defined animal, clothing design, industrial design, and object. This hierarchical labeling enabled precise task construction and increased the benchmark’s diversity.
Task Construction. The editing instructions were generated using Gemini. For each task, we manually selected suitable reference categories, randomly sampled images, and presented them to Gemini to produce corresponding editing instructions. Gemini then evaluated whether following each instruction would lead to visual breakdowns. For instance, in a replacement task, whether the replaced object would appear unnaturally suspended in mid-air.
Among the generated candidates, we retained only the reference–instruction pairs that did not exhibit visual breakdowns and then asked Gemini to further assess whether each instruction constituted a challenging editing scenario, such as cross-domain modifications. Finally, we manually reviewed the remaining data to remove any residual cases that Gemini failed to detect or instructions that were deemed too trivial.
Difficult Reference Categorization. For the cross-domain category, we selected samples whose reference images spanned at least two domains (e.g., color photos, anime, paintings). The different-scale category included samples with both close-up and non–close-up views. The rare-concept category contained samples featuring animals with unusual colors or patterns, humans or animals with atypical makeup or clothing, or fictional creatures. The multilingual category included samples with text in two different languages among English, Chinese, and Japanese. Samples not matching any category were grouped as others.
Table 2: Difficult reference combination statistics and ratios relative to the total 3769 tasks. Our benchmark provides a sufficient number of samples for assessing a model’s compositional reasoning, identity preservation, and cross-domain generalization, compared to prior work [xia2025dreamomni2].
Cross Different Rare Multi-
Domain Scale & View Concept lingual
Count 1063 1357 743 99
Ratio (%) 28.2% 36.0% 19.7% 2.6%
Table 3: Average performance per model across different task categories (10-point scale; the higher the better). The average scores from Gemini-2.5 and GPT-5 are reported. Both open-source and closed-source models exhibit notably lower performance on background replacement tasks, regardless of the number of reference images.
Model Single Two X Objects X-1 + Local X-1 + Global X-1 + Background
Nano Banana 7.817 4.891 4.453 4.118 4.698 3.575
+ Agent (IPR) 7.606 5.030 4.433 4.030 4.496 3.767
GPT-Image-1 7.804 6.585 5.086 5.147 5.757 5.019
+ Agent (IPR) 7.820 6.808 5.258 5.419 5.775 5.284
Qwen-Image-Edit-2509 7.499 3.699 2.256 2.031 2.351 2.033
DreamOmni2 6.520 4.069 2.804 3.037 2.867 2.594
OmniGen2 5.919 3.442 3.256 3.598 3.369 3.022
3.3 Statistics in MultiBanana
Image Category Statistics.   This section presents a statistical overview of both the real and synthetic source data used to construct our benchmark. The dataset is organized into six major categories and thirteen subcategories based on visual and semantic attributes.
In the real dataset, the largest major category is background, comprising 798 images, one-fourth of the total. These mainly consist of images without distinct subjects. The next largest categories are person (741 images) and style (579 images), and together these three categories account for 70% of the dataset. In contrast, light (195 images), object (413 images), and tone (118 images) are relatively small in size. Although these categories are useful as references for changing global conditions such as illumination and overall atmosphere, their sample sizes remain limited.
Within the object category, there are 111 images of animals, 40 of clothing and textile design, 71 of industrial design, and 191 of miscellaneous objects, indicating that data focusing on specific materials or design patterns is scarce. Most images in the background, style, and tone categories lack clear subjects. Furthermore, images containing textual elements, such as logos or signs, are nearly absent. The person category is subdivided into finer attributes: face expression (10), female (181), male (139), hairstyle (135), makeup (59), and pose usefulness (217). While pose-related data account for the largest portion (around 30%), reference data capturing subtle characteristics such as facial expression or makeup remain limited, resulting in a skewed distribution.
To address this imbalance, synthetic data were generated, focusing primarily on images containing explicit subjects (e.g., persons, animals, and objects) as well as those including text. The generation process was designed to intentionally expand the coverage of underrepresented subcategories and improve the overall balance among attributes.
[…]
Consequently, the dataset became more balanced between background-centric and subject-centric images, with broader coverage of diverse visual attributes. This improvement enhances the benchmark’s comprehensiveness and enables more versatile evaluation and instruction-based generation tasks under a wide range of conditions.
Figure 5: Changes in scores for each evaluation criterion when varying the number of reference images. Both open-source and closed-source models exhibit a general trend of decreasing all scores as the number of references increases.
Task Statistics.   After filtering the generated editing instructions using both Gemini and manual checks, we obtained 3,769 high-quality reference images and instruction sets. Among them, 264 sets correspond to single-reference tasks, 907 to two-reference tasks, and 2,598 to multi-reference tasks. Figure 4 provides a further breakdown of the multi-reference tasks, showing that they are evenly distributed with respect to the number of reference images. Additionally, for each task type, X Object, X–1 Object + Background, X–1 Object + Local, and X–1 Object + Global, we ensured that at least 70 sets were included. This is substantially larger than the number of sets in multi-reference tasks of DreamOmni2, which is on the order of ten, enabling more reliable evaluation of model performance. Furthermore, Table 2 shows that our newly introduced evaluation axes – cross-domain, scale and viewpoint differences, rare concept, and multilingual – provide sufficient samples. See Appendix B for details.
Figure 6: Results for diffucult reference combinations. For cross-domain and different scale and view tasks, every model shows lower scores than tasks without such conditions.
3.4 Evaluation Setting
For evaluating the quality of edited images, human evaluation is among the most valuable methods, yet gathering human feedback at scale is prohibitively costly. A practical approach to reduce time and cost is to leverage AI feedback from VLMs (e.g., Gemini-2.5 [geminiteam2023gemini] and GPT-5 [openai2025gpt5]), which has been shown to modestly align with human judgments of image quality [na2024boost, wu2024boosting, furuta2024improving, oshima2025inference]. Therefore, we used Gemini-2.5 and GPT-5 to rate all generated images.
[…]
4 Experiments
We adopt several closed models (i.e., Nano Banana [google2025nanobanana], GPT-Image-1 [openai2025gpt4oimage]) and open models (i.e., Qwen-Image-Edit-2509 [wu2025qwenimagetechnicalreport]], DreamOmni2 [xia2025dreamomni2], OmniGen2 [wu2025omnigen2]) for MultiBanana evaluation. For all models except OmniGen2, we generate outputs for every combination of reference images and text prompts included in the MultiBanana benchmark. Since OmniGen2 does not support more than six reference images, we only evaluate it under the 5-reference setting. As described in Section 3.4, the evaluation is conducted using Gemini-2.5 [geminiteam2023gemini] and GPT-5 [openai2025gpt5]. In the main paper, we report the average scores from Gemini-2.5 and GPT-5, while the individual results for each evaluator are provided in the supplementary material.
Figure 7: (Left) Failure cases of cross-domain, different scale, rare concept, multilingual text. (Right) Failure cases of multi-references in open-source models. They tend to ignore multiple subjects under high-reference conditions.
4.1 Per-Task Evaluation
We compute the MultiBanana scores for each task type defined in Section 3.1, and present the results in Table 3 and Section E.1. For the two-reference tasks, we report the average score across all 11 task variants. For the multi-reference tasks, we compute the average score for each task type (i.e., X Objects, X–1 Objects + Local, X–1 Objects + Global, X–1 Objects + Background) across $X=\{3,4,5,6,7,8\}$.
Overall, both open-source and closed-source models exhibit notably lower performance on background replacement tasks, regardless of the number of reference images. The results for the two-reference setting also reveal clear performance gaps between closed-source and open-source models, even for relatively simple editing scenarios such as single-reference or two-reference tasks.
4.2 Effect of the Number of References
We investigate how the MultiBanana score changes when varying the number of reference images $X=\{3,4,5,6,7,8\}$ for each multi-reference generation task (Figure 5). Both open-source and closed-source models exhibit a general trend of decreasing Total Score as the number of references increases. However, the nature of this degradation differs markedly between the two model families.
Closed-source models demonstrate strong adherence to reference attributes and prompt instructions. As a result, even when the number of references grows, the decline in Instruction Alignment and Reference Consistency remains relatively mild. Nevertheless, because these models attempt to satisfy all reference constraints strictly, increasing the number of subjects often leads to degraded overall visual quality, including over-crowded scenes and compositional inconsistencies (see Figure 1 and Figure 7; left).
In contrast, open-source models show weaker adherence to prompts and references; even under the 3-reference setting, both Instruction Alignment and Reference Consistency remain low, and these scores further deteriorate as the reference count increases. Qualitative inspection also reveals that open models frequently ignore multiple subjects under high-reference conditions. However, their overall quality (background consistency, physical realism, visual quality) remains relatively stable, and they tend not to suffer from the severe compositional collapse observed in closed models (see Figure 7; right).
In summary, closed-source models succeed in incorporating all required subjects but often produce globally inconsistent or distorted scenes due to overfitting to fine-grained reference details. Open-source models, by contrast, tend to omit multiple reference subjects in complex scenarios but generate visually clean images with fewer structural failures. These observations highlight a clear trade-off between strict reference fidelity and holistic image coherence.
4.3 Analysis of Difficult Reference Combinations
Next, to analyze the challenging subject combinations included in this benchmark, namely those described in Section 3.1, we compare the MultiBanana scores between tasks containing these difficult combinations and those that do not in Figure 6. For cross-domain, different-scale and view, rare-concept, and multilingual tasks, most models show lower scores than for tasks without such conditions. These results indicate that the difficult reference concepts focused on by our benchmark remain challenging, even for state-of-the-art closed models. This demonstrates that our benchmark effectively uncovers the current limitations of multi-reference image generation.
4.4 Agentic Framework
To address the performance limitations on our challenging benchmark, we further construct three agentic framework baselines: Iterative Prompt Refinement (IPR), Context-Aware Feedback Generation (CAFG), and Selective Reference Adaptation (SRA). We then conduct experiments on our MultiBanana benchmark. We show the results of IPR in Table 3, where the generator creates images, and based on these images, the planner rewrites the prompt over multiple steps. While GPT’s performance improved with the agent framework, Nano Banana showed no improvement and, in some cases, performance degraded due to information loss from the initial instruction. For details on the problem setting, ablations, and results, see the supplementary material.
5 Conclusion
We constructed a comprehensive benchmark for multi-reference image generation, focusing on challenges specific to this setting, such as difficult reference combinations. Using it, we evaluated state-of-the-art open and closed models and found two notable failure modes as the number of references increases: models either ignore editing instructions or collapse when trying to follow them. Across different types of reference difficulty, tasks involving cross-domain inputs, large-scale or viewpoint gaps, rare concepts, and multilingual inputs are consistently challenging.
Acknowledgements
We thank Minsu Kim and Heiga Zen for their support on this work and review on the initial version of the paper. We also appriciate the funding support from Google Japan. MS was supported by JSPS KAKENHI Grant Number JP23H04974.
[…]
Supplementary Material
A Details on Benchmark Construction
A.1 Details on Image Collection
We constructed the reference image collection for our benchmark using both real images sourced from LAION-5B [laion2022aesthentic] and synthetic images generated by Nano Banana [google2025nanobanana] and GPT-Image-1 [openai2025gpt4oimage]. By combining real and synthetic data, we mitigated the categorical biases inherent in the real-image distribution while substantially expanding the coverage and diversity of the reference set. To generate synthetic images, we designed distinct axes of variation for four major categories: humans, animals, objects, and text-containing images.
Human Category  We systematically combined attributes such as makeup, hairstyle, hair color, pose, strong facial expressions, lighting, and overall tonal style to create a wide range of appearance conditions. Examples include ethnic or fashion-model-style makeup, vivid or unconventional hair colors, model-like or manga-like poses, strong emotional expressions (e.g., anger, joy, surprise), dramatic lighting such as window light through blinds or neon illumination, and global color tones such as sepia or bluish palettes. These conditions were intentionally selected to supplement variations that are underrepresented in real-world data.
Animal Category  We first specified major species groups (dogs, cats, birds, reptiles, wild animals, and fantasy animals). Then we added attributes to enhance the diversity specific to animal imagery. These included human-like behaviors (e.g., a bear reading a book, a fox playing in a rock band), rare colors (e.g., a blue horse or a pink panda), distinctive lighting conditions, and varied global tones such as sepia. This incorporated rare or extreme examples that seldom appear in real datasets.
Object Category  We expanded diversity along two axes: intrinsic object attributes and photographic conditions. The former included distinctive textures (e.g., denim, leather, knit), characteristic prints (e.g., floral, striped, checkered), and variations in material or design (e.g., matte, glossy, transparent). The latter included variations in lighting and global tonal adjustments (e.g., sepia, cool color tones), thereby ensuring a broad range of appearance conditions for objects.
Images Containing Text  Images containing text, which were largely absent from real data, were generated exclusively using GPT-Image-1, as Nano Banana exhibited limited ability to render non-English text. We defined three levels of text length (a single word, a short phrase of up to five words, and a longer sentence). We combined each with two scene types: text rendered directly within a landscape (e.g., ocean, forest) and text appearing naturally in the environment (e.g., on signs in real scenes). These six prompt patterns were generated in English and subsequently translated into Japanese and Chinese to enrich multilingual text-containing images.
Scene Diversity  Finally, all synthetic prompts were designed to incorporate scene diversity, including European cityscapes, university campuses, tropical beaches, event halls, and other varied environments. The generated subjects were ensured to be well-framed, resulting in images suitable for use as high-quality reference material.
A.2 Details on Image Filtering
First, to remove images unsuitable for editing, the images were filtered based on the size of the foreground objects, following the manner in [ye2025imgedit]. YOLOv12 [tian2025yolov12] was applied to all images to detect objects and then performed semantic segmentation using SAM [ravi2025sam], conditioned on the detected bounding boxes. Images were discarded if they met any of the following criteria: the bounding box area covered less than 2% of the entire image, the segmented region within the bounding box covered less than 30%, or the CLIP similarity [ma2022xclip] between the YOLO-predicted class name and the bounding box region was below 20. These cases were judged to contain objects that were not clearly visible. However, images where no objects were detected were retained, assuming they are useful for background or style-level editing tasks. After this filtering, 48% of the images were retained.
We also filtered out images that were inappropriate or unsuitable as references, such as unsafe content, charts, and screenshots of system messages, using both Gemini-based automated screening and human review. For inappropriate content, we specifically targeted categories including hate, harassment, violence, self-harm, sexual content, nudity, shocking content, illegal activity, and other distressing material, ultimately excluding approximately 3% of the collected images. We also manually removed synthetic data that was nearly identical. An example is shown in Figure 8.
Figure 8: Examples of a duplicated synthetic image.
Figure 9: (Left) Number of tasks by reference count. The two-reference tasks contain more samples than the other reference tasks because they include multiple task types. (Right) Breakdown of the two-reference tasks. It contains eleven tasks.
A.3 Details on Category Classification
To construct multi-reference image editing tasks, each reference image had to be accurately categorized. For example, in portrait transformation tasks that modify a person’s hairstyle or makeup, both reference images must contain a person. Similarly, in replacement tasks involving humans or objects, the corresponding entities must be present in the reference images, and for background replacement tasks, background scenes must be determined in the reference images. To satisfy these task-dependent requirements, we performed hierarchical image classification on a large combined set of real and synthetic images.
For synthetic data, attributes can be specified directly during generation, eliminating the need for additional classification. In contrast, real images exhibit substantial variability, requiring precise categorization to determine whether they can serve as valid reference images. To address this, we designed a structured annotation protocol using Gemini that consists of three classification processes for the human, object, and background components of an image.
Human Image Classification  Human classification begins with identifying how many people appear in the image. If there is precisely one person, the model additionally classifies attributes such as gender, makeup, hairstyle distinctiveness, pose clarity, and facial expression characteristics. The output follows a fixed template format, allowing consistent extraction of appearance features relevant to multi-reference generation.
Object Image Classification  Object image classification follows a similar structure. The model first determines whether the image contains no object, a single object, or multiple objects. When exactly one object is detected, it is further categorized according to its visual and material properties, distinguishing among animal-like forms, patterned or decorative textiles, engineered or manufactured materials, and objects lacking distinctive design features.
Background Image Classification  Background classification decomposes the scene into realism, tonal uniformity, and lighting strength. Realism distinguishes photographic scenes from stylized or synthetic ones; tone indicates whether the overall color distribution is uniform or varied; and lighting strength identifies whether a clear directional key light is present. Ambiguous lighting is conservatively labeled as absent to maintain annotation consistency.
A.4 Details on Task Construction
The editing instructions were generated using Gemini. First, we manually selected suitable reference categories for each task, for instance, the person category for the hairstyle modification task. Next, we randomly sampled images from each category and presented them to Gemini to produce corresponding editing instructions. For example, in editing tasks that require object transformations, we first prompted the model to determine which reference image’s object should be modified and which reference image’s attributes (e.g., color) should be used as the target. Afterward, several instruction examples were provided to Gemini, and we asked Gemini to generate new instructions following the same syntax. After that, editing instructions that would cause visual breakdowns were removed by scoring them with Gemini. Additionally, the difficulty of editing instructions was also assessed by Gemini, and those judged easy, such as cases without domain differences, were removed. Finally, the editing instructions were manually verified not to cause image breakdowns or result in overly simple edits, and assigned the remaining samples to the appropriate difficulty categories (cross-domain, scale and viewpoint differences, rare concept, and multilingual text rendering).
B Further Statistics for MultiBanana
The left of Figure 9 shows the number of tasks by each reference count. The two-reference tasks contain more samples than the other tasks because they include multiple task types. The right of Figure 9 shows the breakdown of the two-reference tasks. It contains eleven tasks considered in the prior work [xia2025dreamomni2]: subject addition, subject replacement, background change, color modification, material modification, pose modification, hairstyle modification, makeup modification, tone transformation, style transfer, and text correction. Each task is ensured to include at least 6% of the editing samples, corresponding to roughly 60 samples, which is larger than the number of samples in xia2025dreamomni2.
Figure 10: Qualitative example of “Hard” category of ImgEdit. These tasks are close to being fully solvable by advanced models such as Nano Banana.
Figure 11: Qualitative example of DreamOmni2 benchmark. These tasks are close to being fully solvable by advanced models such as Nano Banana.
C Further Comparison with Prior Benchmarks
Table 1 shows that existing benchmarks do not provide systematic evaluation across diverse multi-reference conditions, support only a limited number of references, and fail to adequately account for important factors such as differences and compatibility among reference images. To illustrate that existing benchmarks are almost solved by state-of-the-art models such as Nano Banana [google2025nanobanana] and GPT-Image-1 [openai2025gpt4oimage], we evaluate the image generation capabilities of these models using ImgEdit [ye2025imgedit], the latest benchmark for image editing, and DreamOmni2 [xia2025dreamomni2], the state-of-the-art benchmark for multi-reference image generation. The quantitative results show that recent closed-source, state-of-the-art models achieve consistently high scores, indicating that these benchmarks are nearing saturation and may soon be unable to meaningfully distinguish among high-performing models (Table 4 and Table 5). The evaluation results are taken from the official GitHub repository of ImgEdit [^1] https://github.com/PKU-YuanGroup/ImgEdit?tab=readme-ov-file and the results reported in the DreamOmni2 paper. Our qualitative results suggest that, even in the additional “Hard” category of ImgEdit, most tasks are already close to being fully solvable by advanced models such as Nano Banana (Figure 10). A similar tendency is observed for DreamOmni2 benchmark as well (Figure 11).
Taken together, these findings indicate that current benchmarks for image editing and multi-reference image generation are approaching their limits in evaluating cutting-edge models. There is a clear need for more challenging benchmarks that can more effectively capture the capabilities and failure modes of the latest generation of image models.
Table 4: Comparison results of different models on ImgEdit-Bench [ye2025imgedit]. “Overall” is computed by averaging the scores across all task types. GPT-4.1 [openai2023gpt4] is used for evaluation. The results show that recent closed-source, state-of-the-art models achieve substantially higher scores, suggesting that the benchmark may be approaching its ceiling in distinguishing high-end models. Evaluation results excluding Nano Banana are taken from the official GitHub repository of ye2025imgedit.
Model Add Adjust Extract Replace Remove Background Style Hybrid Overall
MagicBrush [zhang2023magicbrush] 2.84 1.58 1.51 1.97 1.58 1.75 2.38 1.62 1.83
AnyEdit [yu2025anyedit] 3.18 2.95 1.88 2.47 2.23 2.24 2.85 1.56 2.45
OmniGen2 [wu2025omnigen2] 3.57 3.06 1.77 3.74 3.20 3.57 4.81 2.52 3.44
Kontext-dev [labs2025flux1kontext] 3.83 3.65 2.27 4.45 3.17 3.98 4.55 3.35 3.71
Nano Banana [google2025nanobanana] 3.63 4.66 3.73 4.69 4.71 4.60 4.47 4.10 4.37
GPT-Image-1 [openai2025gpt4oimage] 4.61 4.33 2.90 4.35 3.66 4.57 4.93 3.96 4.20
Table 5: Quantitative comparison of multimodal instruction-based editing and generation in DreamOmni2 Benchmark [xia2025dreamomni2]. Both tasks are evaluated using Gemini [geminiteam2023gemini] and Doubao [bytedance2025doubao]. This result shows that recent closed-source, state-of-the-art models achieve substantially higher scores, indicating that the benchmark may be reaching its ceiling for distinguishing among high-end models. We refer to evaluation results from xia2025dreamomni2.
Method Editing Task Generation Task
Concrete $\uparrow$ Abstract $\downarrow$ Concrete $\uparrow$ Abstract $\downarrow$
Gemini Doubao Gemini Doubao Gemini Doubao Gemini Doubao
Omnigen2 [wu2025omnigen2] 0.2195 0.2927 0.0427 0.0793 0.2083 0.2500 0.1000 0.0778
Kontext [labs2025flux1kontext] 0.0488 0.1220 0.0183 0.0122 0.2500 0.3750 0.0556 0.1222
Qwen-Image-Edit-2509 [wu2025qwenimagetechnicalreport] 0.2683 0.2927 0.0488 0.1159 0.1250 0.2917 0.1111 0.1556
DreamOmni2 [xia2025dreamomni2] 0.5854 0.6585 0.5854 0.6280 0.5833 0.6667 0.5778 0.6333
Nano Banana [google2025nanobanana] 0.6829 0.7073 0.6463 0.5488 0.5000 0.5417 0.5556 0.5488
GPT-Image-1 [openai2025gpt4oimage] 0.6829 0.7805 0.7195 0.7439 0.6250 0.6250 0.6889 0.6333
D Further Discussions for AI and Human Evaluation
The correlation between AI and human evaluation has been shown in prior works [na2024boost, wu2024boosting, furuta2024improving, oshima2025inference]. Consequently, using VLMs is becoming a standard method for the automatic evaluation of generated images [ye2025imgedit, wu2025omnigen2, xia2025dreamomni2]. We used VLMs to evaluate the images along five metrics: Instruction Alignment, Reference Consistency, Background Subject Match, Physical Realism, and Visual Quality. The overall score is calculated as a weighted average of these metrics. The first two metrics represent the fundamental components of reference-based generation, while the last three metrics break down the overall quality. Therefore, the overall score is calculated as $\mathrm{avg}(\mathrm{Instruction\,Alignment}+\mathrm{Reference\,Consistency}+\mathrm{avg}(\mathrm{Background\,Subject\,Match}+\mathrm{Physical\,Realism}+\mathrm{Visual\,Quality}))$.
To demonstrate the consistency of our evaluation metrics, we examined the correlation between different VLMs. We use the evaluated results for Nano Banana and GPT-Image-1, and examine the correlation coefficient between GPT’s and Gemini’s evaluation score. To ensure the reliability of the correlation coefficient, we divided the evaluation results into ten subsets and computed the average correlation across them. Table 6 shows that the evaluated scores exhibit positive correlations, which demonstrates that our evaluation criteria are well-defined and consistent.
Table 6: Correlation coefficients between GPT’s and Gemini’s evaluated scores. $\pm$ shows 95% confidence interval.
Evaluation criteria Correlation coefficients
Instruction Alignment 0.645 ± $\pm\,0.021$
Reference Consistency 0.549 ± $\pm\,0.026$
Background Subject Match 0.601 ± $\pm\,0.020$
Physical Reallism 0.620 ± $\pm\,0.022$
Visual Quality 0.577 ± $\pm\,0.018$
Overall 0.650 ± $\pm\,0.014$
To further ensure the reliability of our evaluation metrics, we examined whether they can detect performance gains from fine-tuning on external image-editing datasets, thereby demonstrating that they align well with those used in other image-editing benchmarks. We compare Qwen-Image-Edit with Qwen-Image-Edit-2509, a fine-tuned variant designed to improve reference consistency and support multi-reference inputs. As shown in Table 7, our evaluation metrics successfully capture the performance gains of Qwen-Image-Edit-2509 over Qwen-Image-Edit. We conducted this comparison using single-reference tasks because Qwen-Image-Edit supports only a single input image.
Table 7: Comparison of Qwen-Image-Edit and Qwen-Image-Edit-2509 on single-reference task.
Qwen-Image-Edit Qwen-Image-Edit-2509
single 6.332 7.499
E Detailed Results
E.1 Further Discussion of Per-Task Evaluation
Figure 12, Figure 13, and Figure 14 show the scores of each model across five evaluation metrics for each task, evaluated by GPT, Gemini, and their average, respectively. Taking the weighted average across five metrics, Figure 15 shows the overall score of each model for single and two-reference tasks, evaluated by GPT, Gemini, and their average, respectively. For multi-reference tasks, Table 8, Table 9, and Table 10 show the overall score, evaluated by GPT, Gemini, and their average, respectively. Overall, we observed a substantial gap in Instruction Alignment and Reference Consistency between the closed-source models and the other open-source models. In particular, GPT-Image-1 achieves significantly superior performance in Instruction Alignment, especially in style and text modification tasks in the two-reference task. Meanwhile, in Reference Consistency, Nano Banana achieves the highest scores in the single-reference task. Nano Banana also performs on par with GPT-Image-1 in Reference Consistency, across background modification, subject replacement, subject addition, and makeup tasks, which indicates its strong ability to leverage reference images, particularly when the number of references is small. Additionally, Qwen-Image-Edit-2509, the most advanced open-source model, performs worse than the other models in background modification tasks in Background Subject Match and Visual Quality. This result suggests that it struggles with background modification.
Figure 12: Scores of each model across evaluation metrics for each task by GPT. The horizontal axis denotes the scores for the five evaluation criteria, and the vertical axis denotes the number of reference images.
Figure 13: Scores of each model across evaluation metrics for each task by Gemini. The horizontal axis denotes the scores for the five evaluation criteria, and the vertical axis denotes the number of reference images.
Figure 14: Average scores of each model across evaluation metrics for each task by GPT and Gemini. The horizontal axis denotes the scores for the five evaluation criteria, and the vertical axis denotes the number of reference images.
Figure 15: Overall scores of each model for single and two-reference tasks by GPT, Gemini, and their average.
Table 8: Detailed per-task overall scores for the multi-reference tasks, evaluated by GPT. The local, global, back, and object columns under X correspond to X–1 Objects + Local, X–1 Objects + Global, X–1 Objects + Background, and X Object, respectively.
Model 3-references 4-references 5-references
local global back object local global back object local global back object
DreamOmni2 3.94 4.19 3.92 3.91 3.34 3.24 2.87 3.20 3.30 3.06 2.73 3.22
OmniGen2 4.42 4.59 4.47 4.47 4.41 3.37 3.10 3.27 3.78 3.28 2.87 3.31
Qwen-Image-Edit-2509 4.42 4.65 4.75 4.77 3.08 3.35 2.78 3.54 1.93 2.40 1.98 2.10
Nano Banana 5.44 5.81 5.10 5.64 4.57 5.36 4.27 5.25 5.12 5.93 3.96 4.93
GPT-image-1 6.58 7.40 6.98 6.40 6.18 6.98 6.73 6.56 6.18 6.66 6.22 6.28
Model 6-references 7-references 8-references
local global back object local global back object local global back object
DreamOmni2 3.37 3.03 2.82 2.96 3.24 3.02 2.71 2.86 3.26 2.79 2.57 2.82
OmniGen2 - - - - - - - - - - - -
Qwen-Image-Edit-2509 1.85 2.18 1.68 1.79 1.51 1.89 1.57 1.88 1.68 2.29 1.62 1.83
Nano Banana 4.43 4.88 3.89 4.90 4.40 4.78 3.57 4.81 4.31 4.82 3.21 4.45
GPT-image-1 6.16 6.08 5.68 5.90 5.38 6.09 5.59 5.94 4.96 5.82 5.03 5.15
Table 9: Detailed per-task overall scores for the multi-reference tasks, evaluated by Gemini. The local, global, back, and object columns under X correspond to X-1 Objects + Local, X–1 Objects + Global, X–1 Objects + Background, and X Object, respectively.
Model 3-references 4-references 5-references
local global back object local global back object local global back object
DreamOmni2 3.00 2.49 2.45 2.74 2.84 2.73 2.38 2.44 2.63 2.50 2.36 2.66
OmniGen2 3.03 2.89 3.00 3.25 3.10 3.06 2.27 2.73 2.84 3.02 2.41 2.50
Qwen-Image-Edit-2509 3.27 3.16 3.29 3.97 2.29 3.00 2.15 2.79 1.14 1.69 1.21 1.20
Nano Banana 4.34 4.69 4.61 4.89 3.42 4.34 3.79 4.14 4.25 4.50 3.16 3.85
GPT-image-1 5.48 6.28 5.37 5.16 4.83 5.26 4.75 5.04 4.89 5.20 4.25 4.10
Model 6-references 7-references 8-references
local global back object local global back object local global back object
DreamOmni2 2.47 2.54 2.14 2.36 2.65 2.50 2.15 2.30 2.42 2.33 2.04 2.19
OmniGen2 - - - - - - - - - - - -
Qwen-Image-Edit-2509 1.00 1.26 1.00 1.03 1.01 1.06 1.33 1.17 1.18 1.31 1.01 1.00
Nano Banana 3.25 3.91 2.55 3.79 3.05 3.73 2.26 3.52 2.85 3.64 2.53 3.25
GPT-image-1 3.66 4.66 3.46 4.02 4.18 4.58 3.24 3.52 3.28 4.07 2.91 2.96
Table 10: Detailed per-task overall scores for each multi-reference task, evaluated by the average of GPT and Gemini. The local, global, back, and object columns under X correspond to X–1 Objects + Local, X–1 Objects + Global, X–1 Objects + Background, and X Object, respectively.
Model 3-references 4-references 5-references
local global back object local global back object local global back object
DreamOmni2 3.47 3.34 3.18 3.33 3.09 2.98 2.62 2.82 2.97 2.78 2.54 2.94
OmniGen2 3.73 3.74 3.74 3.86 3.76 3.22 2.69 3.00 3.31 3.15 2.64 2.91
Qwen-Image-Edit-2509 3.85 3.90 4.02 4.37 2.69 3.17 2.47 3.17 1.54 2.04 1.60 1.65
Nano Banana 4.89 5.25 4.85 5.27 4.00 4.85 4.03 4.70 4.69 5.22 3.56 4.39
GPT-image-1 6.03 6.84 6.18 5.78 5.50 6.12 5.74 5.80 5.54 5.93 5.24 5.19
Model 6-references 7-references 8-references
local global back object local global back object local global back object
DreamOmni2 2.92 2.78 2.48 2.66 2.94 2.76 2.43 2.58 2.84 2.56 2.31 2.50
OmniGen2 - - - - - - - - - - - -
Qwen-Image-Edit-2509 1.42 1.72 1.34 1.41 1.26 1.47 1.45 1.52 1.43 1.80 1.32 1.42
Nano Banana 3.84 4.39 3.22 4.35 3.72 4.25 2.92 4.16 3.58 4.23 2.87 3.85
GPT-image-1 4.91 5.37 4.57 4.96 4.78 5.33 4.42 4.73 4.12 4.94 3.97 4.05
Figure 16: Overall scores for each difficult reference combination across models. Darker colors represent the average score of tasks that include the corresponding combination, while lighter colors indicate the average score of tasks that do not include it.
Figure 17: Qualitative example for 4-Object Tasks.
Figure 18: Qualitative example for 3-Object + Background Tasks.
Figure 19: Qualitative example for 3-Object + Local Tasks.
Figure 20: Qualitative example for 3-Object + Global Tasks.
Figure 21: Qualitative example for Tasks with 8 references.
E.2 Further Discussion of the Effect of the Number of References
According to the per-metric scores in Figure 14 and the overall scores in Table 10, all models exhibit decreasing performance as the number of reference images increases. For Instruction Alignment and Reference Consistency, GPT-Image-1 and Nano Banana achieve relatively high scores, whereas the open-source model, DreamOmni2, tends to approach the minimum score of 1 as the number of references increases. In contrast, for quality-related metrics, Background Subject Match, Physical Realism, and Visual Quality, the DreamOmni2 maintains high scores even when more reference images are provided. As discussed in Section 4, this is because closed-source models prioritize adherence to the references and editing instructions over visual quality, while open-source models show a tendency to preserve visual quality but often ignore the references.
In the finer-grained task-level evaluation, we found that even as the number of reference images increased, the X–1 Object + Global task did not exhibit score degradation in Background Subject Match, Physical Realism, or Visual Quality. This is because the task requires modifying the image to a specified style; therefore, even if the reference images come from different domains, which may degrade the visual quality, the final output converges to a unified style.
In the results of Qwen-Image-Edit-2509, the model is trained to handle up to four reference images; therefore, when provided with five or more references, the input becomes out-of-distribution. This causes the model to produce perceptually invalid outputs, and then the scores approach the minimum value of 1.
E.3 Further Discussion of Difficult Reference Combination
Figure 16 shows the scores of each model in difficult reference combination tasks across evaluation metrics.
Cross-domain diversity   In cross-domain cases, samples with reference images from different domains (darker color) were evaluated as lower Reference Consistency scores, because all models aim to unify the domains during generation.
Scale and viewpoint differences   In the different scale and viewpoint cases, the Reference Consistency scores declined consistently across all models. According to the reasoning traces of the VLM, the primary reasons for deductions are the loss of fine details and changes in pose. This suggests that when models attempt to reproduce objects under different scales or viewpoints, they may fail to preserve fine details or may adjust the pose to achieve a more natural appearance.
Rare concept references   In the rare concept case, because the reference images often contain subjects at a large scale, instances were observed in which the model reproduced this scale without adjustment, even when it was too large. For this reason, rare concept samples tend to receive lower scores on the Physical Realism.
Multilingual references   In the multilingual text rendering case, we observed a consistent decline across all metrics except Reference Consistency. This primarily occurs because models have limited text-rendering capability in languages other than English, leading to failures in cross-language conversion. As a result, instruction alignment and visual quality become lower. However, scores dropping in GPT-Image-1 were lower than those of others, since it exhibits comparatively stronger text-rendering ability in non-English languages.
E.4 Agentic Inference
Figure 22: Example of changes in the Iterative Prompt Refinement (IPR) framework. Comparison and evaluation results between Nano Banana and GPT on the task of generating images from 8 object references according to the instruction prompt.
We introduced three agentic frameworks, Iterative Prompt Refinement (IPR), Context-Aware Feedback Generation (CAFG), and Selective Reference Adaptation (SRA). Here, we refer to $\mathrm{Gen}$ as the Generator, which produces an image, and $\mathrm{Plan}$ as the Planner, which updates the instruction prompt (and selects reference images) for the next step. Let $P^{t}$ denote the instruction prompt at step $t$, $\{R_{i}\}_{i\in[I]}$ the reference images, and $G^{t}$ the generated image with $G^{0}=\varnothing$. We conducted experiments with both GPT and Gemini. For GPT, we use GPT-Image-1 as the Generator and GPT-5 as the Planner. For Gemini, we use Nano Banana as the Generator and Gemini-2.5-Flash as the Planner.
E.4.1 Iterative Prompt Refinement (IPR)
IPR framework is formulated as follows. The Planner refines the prompt based on the generation result from the previous step.
$G^{t+1}=\mathrm{Gen}(P^{t},\{R_{i}\}_{i\in[I]},\varnothing),$
$P^{t+1}=\mathrm{Plan}(P^{t},\{R_{i}\}_{i\in[I]},G^{t+1}).$
E.4.2 Context-Aware Feedback Generation (CAFG)
CAFG framework is formulated as follows. The Generator generates the image based on the generation result (context) from the previous step.
$G^{t+1}=\mathrm{Gen}(P^{t},\{R_{i}\}_{i\in[I]},\bm{G^{t}}),$
$P^{t+1}=\mathrm{Plan}(P^{t},\{R_{i}\}_{i\in[I]},G^{t+1}).$
E.4.3 Selective Reference Adaptation (SRA)
SRA framework is formulated as follows. The Planner selects only the reference images that should be improved based on the generation result from the previous step, and the Generator receives these as context. SRA is expected to reduce task complexity for the agent by adaptively decreasing the number of references.
$G^{t+1}=\mathrm{Gen}(P^{t},\{R_{i}\}_{\bm{i\in U^{t}}},G^{t}),$
$P^{t+1},\bm{U^{t+1}}=\mathrm{Plan}(P^{t},\{R_{i}\}_{i\in[I]},G^{t+1}).$
where $U^{t}$ is the index set of reference images that are insufficiently reflected in the generated image $G^{t}$.
E.4.4 Experiments
We set the maximum number of steps $t$ as $3$ and we evaluate three agentic frameworks—IPR, CAFG, and SRA—using Gemini, and the IPR framework using GPT. Figure 22 illustrates an example of step-wise improvement in the IPR framework. Figure 23, Figure 24, Figure 25, and Figure 26 present the evaluation results of multi-reference generations at each step, averaged across judgments from Gemini-2.5-Flash and GPT-5, broken down by category. Figure 27 presents the results for single and two-reference generations. Nano Banan (Gemini) shows modest improvements in Physical Realism and Visual Quality across refinement steps, while Instruction Alignment and Reference Consistency either remain unchanged or deteriorate. In contrast, GPT demonstrates consistent improvements across all categories. This suggests that Gemini’s planner progressively loses information from the original prompt as refinement steps proceed.
IPR Framework with Gemini
Figure 23: Detailed results of multi-reference image generation using the IPR framework with Gemini.
CAFG Framework with Gemini
Figure 24: Detailed results of multi-reference image generation using the CAFG framework with Gemini.
SRA Framework with Gemini
Figure 25: Detailed results of multi-reference image generation using the SRA framework with Gemini.
IPR Framework with GPT
Figure 26: Detailed results of multi-reference image generation using the IPR framework with GPT.
Figure 27: Detailed results of single and two-reference image generation.
F Extended Related Works
F.1 Controllable Text-to-Image Generation
Diffusion models have achieved remarkable progress in generative modeling, demonstrating exceptional performance, particularly in high-fidelity image synthesis. Representative systems, such as Stable Diffusion [rombach2022ldm, podell2024sdxl, esser2024sd3], FLUX [flux2024, labs2025flux1kontext], DALL-E [ramesh2022dalle], and Imagen [saharia2022imagen, imagenteamgoogle2024imagen3], have demonstrated strong text-to-image generation capabilities, establishing a controllable and scalable foundation for image generation. To enhance controllability, models such as ControlNet [zhang2023controlnet] and T2I-Adapter [mou2023t2i-adapter] introduced external conditioning modules, enabling image-conditioned generation. Recent large pretrained diffusion models with textual conditioning, such as Stable Diffusion [rombach2022ldm, podell2024sdxl, esser2024sd3] and FLUX [flux2024, labs2025flux1kontext], enable the synthesis of novel scene images while preserving the identity depicted in the reference images. Building on these developments, approaches based on fine-tuning (e.g., DreamBooth [ruiz2022dreambooth]) and adapter-based methods (e.g., IP-Adapter [ye2023ip-adapter]) have been actively explored to address this challenge [jiang2025infiniteyou, gal2023an, brooks2023instructpix2pix, mokady2023null, tan2025ominicontrol, NEURIPS2023_602e1a5d]. Additionally, training-free approaches (e.g., pix2pix-zero [parmar2023zero] and prompt-to-prompt [hertz2023prompttoprompt]) have also been widely proposed [Cao_2023_ICCV, Huberman-Spiegelglas_2024_CVPR, Wallace_2023_CVPR, Tumanyan_2023_CVPR, kulikov2025flowedit, miyake2025negative-prompt-inversion]. These advancements collectively demonstrate that diffusion models are becoming increasingly flexible and adaptable across diverse conditional generation settings.
In recent years, research on unified image generation has gained momentum, aiming to handle multiple image generation tasks within a single model. For example, ChatGPT-Image-1 [openai2025gpt4oimage] and Nano Banana [google2025nanobanana] are capable of jointly processing text and image inputs, enabling integrated image generation and editing within a unified framework. Similarly, the open-source models, such as Qwen-Image [wu2025qwenimagetechnicalreport] and FLUX.1-Kontext-dev [labs2025flux1kontext], built upon a robust multimodal large language model backbone, demonstrate high-quality and flexible image generation and editing performance. Beyond these, numerous open-source unified generative models continue to emerge [xiao2024omnigen, xia2025dreamomni, wu2025omnigen2, xia2025dreamomni2].
With the advent of multimodal foundation models capable of jointly handling text and images [google2025nanobanana, openai2025gpt4oimage], a new task has recently become feasible: multi-reference-based image generation. In this task, the model receives multiple reference images along with textual instructions and generates a new scene while preserving the identity of the reference images. However, despite its novelty and potential, a reliable benchmark for evaluating this innovative task has yet to be established, leaving it as an essential open challenge for future research.
F.2 Benchmarks for Reference-Based Generation
Among reference-driven image generation tasks, the most widely recognized one is image editing. Image editing refers to the task of generating an edited image based on a single reference image and a textual instruction describing how the edit should be performed. Representative benchmarks such as MagicBrush [zhang2023magicbrush] and EMU-Edit [sheynin2024emuedit] have extended task-specific evaluations but still rely heavily on similarity-based metrics. SmartEdit [huang2024smartedit] supports editing of complex scenes, but does not sufficiently cover more general settings. Similarly, I2E-Bench [ma2024i2ebench] employs GPT-4o to provide human-aligned evaluations across a variety of editing tasks, yet it uses different metrics for each task and therefore fails to capture the shared characteristics of editing as a whole. More recently, ImgEdit [ye2025imgedit] has enabled the evaluation of multi-turn editing tasks, where image generation and editing are alternated interactively.
On the other hand, one of the earliest benchmarks for reference-based image generation beyond editing is the DreamBooth [ruiz2022dreambooth] dataset. However, such benchmarks are still limited to generation tasks conditioned on a single reference image and thus cannot handle more complex or compositional settings.
In recent years, following advances demonstrated by GPT-Image-1 [openai2025gpt4oimage] and Nano Banana [google2025nanobanana], increasing attention has been directed toward multi-reference image generation, where multiple reference images are used jointly for generation. Benchmarks such as OmniContext [wu2025omnigen2] and DreamOmni2 [xia2025dreamomni2] have been proposed to evaluate this task. Nevertheless, these benchmarks encompass only a small number of images and task types, thereby limiting the scope of the evaluation. Moreover, despite being designed for multi-reference generation, they do not evaluate qualitative differences or relationships between reference images, and their setups remain similar to those of conventional single-image editing benchmarks.
G Prompts
G.1 Prompts of Agentic Framework
Prompt of the Generator in the IPR framework {prompt}{reference_image_files}
Prompt of the Planner in the IPR framework You are a prompt-refiner agent. Previous prompt: {previous_prompt} You are given multiple reference images and one generated image. The generated image is the last one, and the reference images come before it. Based on the reference images and the generated image (the last image), propose a refined prompt that improves how reference images are used and adjusts the composition/style. Return only the new prompt text. {reference_image_files}{previously_generated_image_file}
Prompt of the Generator in the CAFG framework {prompt} The last image provided is the previously generated image from the last step and all images before it are reference images. Please refine the previous generation using the reference images and enhance composition/style. {reference_image_files}{previously_generated_image_file}
Prompt of the Planner in the CAFG framework You are a prompt-refiner agent. Previous prompt: {previous_prompt} You are given multiple reference images and one generated image. The generated image is the last one, and the reference images come before it. Based on the reference images and the generated image (the last image), propose a refined prompt that improves how reference images are used and adjusts the composition/style. Return only the new prompt text. {reference_image_files}{previously_generated_image_file}
Prompt of the Generator in the SRA framework {prompt} The last image provided is the previously generated image from the last step and all images before it are reference images. Please refine the previous generation using the reference images and enhance composition/style. {reference_image_files}{previously_generated_image_file}
Prompt of the Planner in the SRA framework You are a prompt-refiner and reference-selector agent. Previous prompt: {previous_prompt} You are given multiple reference images (in order) and one generated image (the last one). Your task: 1. Identify which reference images are insufficiently reflected in the generated image. 2. Propose a refined prompt that improves the generation. 3. Return your response in JSON format ONLY. Do not include any other text. JSON format: { ”indices”: [0, 2, 3], ”prompt”: ”Your refined prompt text here” } Where ’indices’ is an array of 0-based indices (from 0 to {len(reference_image_files)-1}) of reference images that are insufficiently reflected in the generated image. ’prompt’ is the refined instruction prompt for the next generation step. Example response: { ”indices”: [0, 2, 3], ”prompt”: ”Focus more on the lighting from the first image and the composition from images 3 and 4. Emphasize the color palette and texture details.” } {reference_image_files}{previously_generated_image_file}
G.2 Prompts of VLM as Judge
[…]

### v2

@@
 Abstract
-Recent text-to-image generation models have acquired the ability of multi-reference generation and editing; the ability to inherit the appearance of subjects from multiple reference images and re-render them under new contexts.
-However, the existing benchmark datasets often focus on the generation with single or a few reference images, which prevents us from measuring the progress on how model performance advances or pointing out their weaknesses, under different multi-reference conditions.
-In addition, their task definitions are still vague, typically limited to axes such as “what to edit” or “how many references are given”, and therefore fail to capture the intrinsic difficulty of multi-reference settings.
-To address this gap, we introduce MultiBanana, which is carefully designed to assesses the edge of model capabilities by widely covering multi-reference-specific problems at scale: (1) varying the number of references, (2) domain mismatch among references (e.g., photo vs.
+Recent text-to-image generation models have acquired the ability of multi-reference generation and editing; that is, to inherit the appearance of subjects from multiple reference images and re-render them in new contexts.
+However, existing benchmark datasets often focus on generation using a single or a few reference images, which prevents us from measuring progress in model performance or identifying weaknesses when following instructions with a larger number of references.
+In addition, their task definitions are still vague, limited to axes such as “what to edit” or “how many references are given”, and therefore fail to capture the challenges inherent in combining heterogeneous references.
+To address this gap, we introduce MultiBanana, which is designed to assess the edge of model capabilities by widely covering problems specific to multi-reference settings: (1) varying the number of references (up to 8), (2) domain mismatch among references (e.g., photo vs.
 anime), (3) scale mismatch between reference and target scenes, (4) references containing rare concepts (e.g., a red banana), and (5) multilingual textual references for rendering.
-Our analysis among a variety of text-to-image models reveals their superior performances, typical failure modes, and areas for improvement.
-MultiBanana will be released as an open benchmark to push the boundaries and establish a standardized basis for fair comparison in multi-reference image generation.
+Our analysis among a variety of text-to-image models reveals their respective performances, typical failure modes, and areas for improvement.
+MultiBanana is released as an open benchmark to push the boundaries and establish a standardized basis for fair comparison in multi-reference image generation.
 Our data and code are available at https://github.com/matsuolab/multibanana.
 Figure 1: The overview of MultiBanana.
-MultiBanana widely covers multi-reference specific problems, varying the number of references (the top row), domain and scale mismatch among references (two on the left in the middle row), multilingual text rendering (center in the bottom row), and containing rare concepts (right in the bottom row).
+MultiBanana broadly covers problems specific to multi-reference settings, including varying the number of references (top row), domain and scale mismatches among references (two on the left in the middle row), multilingual text rendering (center in the bottom row), and the presence of rare concepts (right in the bottom row).
 […]
 1 Introduction
-Recent advances in image generation, driven by large-scale multimodal LLM backbones, have enabled unprecedented levels of instruction-following [google2025nanobanana, openai2025gpt4oimage, wu2025qwenimagetechnicalreport, cao2025hunyuanimage3, wu2025less, xiao2024omnigen, wu2025omnigen2, xia2025dreamomni, xia2025dreamomni2, cui2025emu35nativemultimodalmodels].
-Among these developments, multi-reference image generation has emerged as a particularly notable capability.
+Recent advances in image generation, driven by large-scale multimodal LLM backbones, have enabled unprecedented levels of instruction-following [22, 56, 79, 7, 81, 85, 80, 84, 83, 13].
+Among these developments, multi-reference image generation has emerged as a notable capability.
 When a user provides multiple reference images, the model can inherit the subject’s appearance and re-render it in a new context while faithfully preserving identity.
-This shift moves image generation beyond single-reference conditioning toward greater controllability, driving increasing demand in industrial domains such as content production [ruiz2022dreambooth, ye2023ip-adapter], advertising [horita2024retrievalaugmented, morita2025tkg], and fashion design [zhu2023tryondiffusion, choi2024improving, fang2024vivid], where reference-guided generation is essential.
-However, despite this progress, we lack a benchmark that systematically evaluates how model performance varies under different multi-reference conditions.
-As shown in Table 1, existing benchmarks [wang2023editbench, basu2023editval, sheynin2024emuedit, zhang2023magicbrush, huang2024smartedit, yu2025anyedit, ma2024i2ebench, ye2025imgedit, ruiz2022dreambooth, wu2025omnigen2, xia2025dreamomni2] suffer from two fundamental limitations.
-First, they have an extremely narrow coverage due to the small number of unique reference images, making it impossible to assess the impact of varying reference conditions in both single- and multi-reference tasks.
-Second, their task definitions still lack details, typically limited to axes such as“what to edit” or “how many references are given”, and therefore fail to capture the intrinsic difficulty of multi-reference settings.
-Moreover, existing benchmarks limit the number of references to a narrow range (typically 1–5) and do not examine how performance changes with more references.
-To address these limitations, we introduce a new evaluation framework grounded not only in task categories but also in the intrinsic properties of the reference set itself.
+This shift moves image generation beyond single-reference conditioning toward greater controllability, driving increasing demand in industrial domains such as content production [66, 90, 78], advertising [33, 28, 51], and fashion design [98, 11, 16], where reference-guided generation is essential.
+Despite this progress, we lack a benchmark to evaluate how well models follow instructions that involve multiple references under varying conditions.
+As shown in Table 1, existing benchmarks [77, 3, 69, 93, 30, 92, 47, 91, 66, 80, 83] suffer from two fundamental limitations.
+First, existing benchmarks limit the number of references to a narrow range (typically 1–4) and do not examine how well models perform when following instructions with a larger number of references.
+Second, their task definitions are still vague, limited to axes such as “what to edit” or “how many references are given”, and therefore fail to capture the challenges inherent in combining heterogeneous references.
+To address these limitations, we introduce a new evaluation framework that is grounded in the core challenges of multi-reference generation.
 Our benchmark incorporates multiple orthogonal axes specific to the multi-reference regime, including broader coverage of the number of references (up to 8), domain gaps between references (e.g., photo–anime mixtures), mismatches in object scale between reference and target scenes, rare concepts in the references (e.g., a red banana), and multilingual textual rendering references (English, Chinese, and Japanese).
-Using this benchmark, we conduct a cross-conditional analysis to assess model capabilities, identify characteristic failure modes, and point out remaining gaps.
-For example, while performance degradation due to increasing reference counts can be mitigated by agent-style multi-stage generation pipelines, failures arising from unsupported styles or languages remain difficult to remedy with inference-time strategies.
-We will open-source MultiBanana to establish a standard foundation for fair comparison and to accelerate progress in multi-reference image generation.
+Using this benchmark, we analyze model capabilities, identify characteristic failure modes, and highlight remaining gaps.
+For example, our analysis reveals that as instruction difficulty increases (through more references, cross-domain mixtures, or rare concepts), closed-source models (e.g., Nano Banana) tend to incorporate all required subjects but suffer from compositional distortion due to overfitting to reference details, while open-source models (e.g., Qwen-Image) generate visually cleaner outputs but often omit some reference subjects altogether.
+We release MultiBanana as an open-source benchmark to establish a standard foundation for fair comparison and to accelerate progress in multi-reference image generation.
 Table 1: Comparison among major benchmarks for reference-based image generation.
@@
 Benchmark #Size #Tasks #References Difficult Reference Combination Metrics
-EditBench [wang2023editbench] 240 1 1 CLIP [radford2021clip]
-EditVal [basu2023editval] 648 13 1 CLIP, VLM, manual
-EmuEdit [sheynin2024emuedit] 3055 7 1 L1, CLIP, DINO [caron2021dino]
-MagicBrush [zhang2023magicbrush] 1053 9 1 L1, L2, CLIP, DINO
-AnyEdit [yu2025anyedit] 1250 25 1 L1, CLIP, DINO
-I2EBench [ma2024i2ebench] 2240 16 1 GPT [openai2023gpt4]
-ImgEdit-Bench [ye2025imgedit] 811 14 1 GPT (3 dim.), Fake Det. [xu2024fakeshield]
-DreamBooth [ruiz2022dreambooth] 75 1 1 CLIP. DINO
-OmniContext [wu2025omnigen2] 400 8 3 GPT (3 dim.)
-DreamOmni2 [xia2025dreamomni2] 319 29 4 Gemini [geminiteam2023gemini], Doubao [bytedance2025doubao]
+EditBench [77] 240 1 1 CLIP [62]
+EditVal [3] 648 13 1 CLIP, VLM, manual
+EmuEdit [69] 3055 7 1 L1, CLIP, DINO [8]
+MagicBrush [93] 1053 9 1 L1, L2, CLIP, DINO
+AnyEdit [92] 1250 25 1 L1, CLIP, DINO
+I2EBench [47] 2240 16 1 GPT [54]
+ImgEdit-Bench [91] 811 14 1 GPT (3 dim.), Fake Det. [87]
+DreamBooth [66] 75 1 1 CLIP. DINO
+OmniContext [80] 400 8 3 GPT (3 dim.)
+DreamOmni2 [83] 319 29 4 Gemini [19], Doubao [5]
 MultiBanana (Ours) 3769 36 8 GPT, Gemini (5 dim.)
@@
 2.1 Controllable Text-to-Image Generation
-Diffusion models have achieved state-of-the-art performance in high-fidelity image synthesis.
-Recent large pretrained diffusion models with textual conditioning, such as Stable Diffusion [rombach2022ldm, podell2024sdxl, esser2024sd3] and FLUX [flux2024, labs2025flux1kontext], enable the synthesis of novel scene images while preserving the identity depicted in the reference images.
-Building on these developments, approaches based on fine-tuning (e.g., DreamBooth [ruiz2022dreambooth]) and adapter-based methods (e.g., IP-Adapter [ye2023ip-adapter]) have been actively explored to address the controllability [jiang2025infiniteyou, gal2023an, brooks2023instructpix2pix, mokady2023null, tan2025ominicontrol, NEURIPS2023_602e1a5d].
-Additionally, training-free approaches (e.g., prompt-to-prompt [hertz2023prompttoprompt]) have also been widely proposed [parmar2023zero, Cao_2023_ICCV, Huberman-Spiegelglas_2024_CVPR, Wallace_2023_CVPR, Tumanyan_2023_CVPR, kulikov2025flowedit, miyake2025negative-prompt-inversion].
-Recent research on unified image generation has gained momentum, aiming to handle multiple image generation tasks within a single model.
-For example, ChatGPT-Image-1 [openai2025gpt4oimage] and Nano Banana [google2025nanobanana] can jointly process text and image inputs, enabling integrated image generation and editing within a unified framework.
-Similarly, the open-source models, such as Qwen-Image [wu2025qwenimagetechnicalreport] and FLUX.1-Kontext-dev [labs2025flux1kontext], built upon a robust multimodal large language model backbone, demonstrates high-quality and flexible image generation and editing performance.
-Beyond these, numerous open-source unified generative models continue to emerge [xiao2024omnigen, xia2025dreamomni, wu2025omnigen2, xia2025dreamomni2].
-These advancements collectively demonstrate that diffusion models are becoming increasingly flexible and adaptable across diverse conditional generation settings.
+Diffusion models have achieved state-of-the-art performance in high-fidelity image synthesis [70, 26].
+Large pretrained diffusion models with textual conditioning, such as Stable Diffusion [65, 60, 15] and FLUX [42, 41], can generate novel images while preserving the identity depicted in the reference images.
+Building on these developments, approaches based on fine-tuning (e.g., DreamBooth [66]) and adapter-based methods (e.g., IP-Adapter [90]) have been explored to improve the controllability [34, 18, 4, 50, 72, 44].
+Recent research on multimodal LLM-based image generation has gained momentum, aiming to handle multiple image generation tasks within a single model.
+For example, GPT-Image-1 [56] and Nano Banana [22] can jointly process text and image inputs, enabling integrated image generation and editing within a unified framework.
+Similarly, open-source models such as Qwen-Image [79] and FLUX.1-Kontext-dev [41] demonstrate high-quality, flexible image generation and editing.
+Beyond these, numerous open-source unified generative models continue to emerge [85, 84, 80, 83].
+These advancements demonstrate that diffusion models are becoming flexible and adaptable across diverse conditional generation settings.
 2.2 Benchmarks for Reference-Based Generation
-Among reference-driven image generation tasks, image editing can be the most widely recognized, which has referred to the task of generating an edited image based on a single reference image and a textual instruction describing how the edit should be performed.
-Representative benchmarks such as MagicBrush [zhang2023magicbrush] and EMU-Edit [sheynin2024emuedit] have extended task-specific evaluations but still rely heavily on similarity-based metrics.
-SmartEdit [huang2024smartedit] supports editing of complex scenes, but does not sufficiently cover more general settings.
-Similarly, I2E-Bench [ma2024i2ebench] employs GPT-4o to provide human-aligned evaluations across a variety of editing tasks, yet it uses different metrics for each task and therefore fails to capture the shared characteristics of editing as a whole.
-More recently, ImgEdit [ye2025imgedit] has enabled the evaluation of multi-turn editing tasks, where image generation and editing are alternated interactively.
-One of the earliest benchmarks for reference-based image generation beyond instruction-based editing is the DreamBooth [ruiz2022dreambooth] dataset.
+Among reference-driven image generation tasks, image editing is the most widely recognized: it involves generating an edited image from a single reference image and a textual instruction describing how the edit should be performed.
+Representative benchmarks such as MagicBrush [93] and EMU-Edit [69] have extended task-specific evaluations but still rely heavily on similarity-based metrics.
+SmartEdit [30] supports editing of complex scenes, but does not sufficiently cover more general settings.
+Similarly, I2E-Bench [47] employs GPT-4o to provide human-aligned evaluations across a variety of editing tasks, yet it uses different metrics for each task and therefore fails to capture the shared characteristics of editing as a whole.
+ImgEdit [91] has enabled evaluation of multi-turn editing tasks, in which image generation and editing are alternated interactively.
+One of the earliest benchmarks for reference-based image generation beyond instruction-based editing is the DreamBooth [66] dataset.
 However, such benchmarks are still limited to generation tasks conditioned on a single reference image and thus cannot handle more complex or compositional settings.
-Recently, following advances demonstrated by GPT-Image-1 [openai2025gpt4oimage] and Nano Banana [google2025nanobanana], increasing attention has been directed toward multi-reference image generation, where multiple reference images are used jointly for generation.
-Benchmarks such as OmniContext [wu2025omnigen2] and DreamOmni2 [xia2025dreamomni2] have been proposed to evaluate this task.
+Recently, advances demonstrated by GPT-Image-1 [56] and Nano Banana [22] have drawn increasing attention to multi-reference image generation, where multiple reference images are jointly used for generation.
+Benchmarks such as OmniContext [80] and DreamOmni2 [83] have been proposed to evaluate this task.
 Nevertheless, these benchmarks encompass only a small number of images and task types, thereby limiting the scope of the evaluation.
-Moreover, despite being designed for multi-reference generation, they do not evaluate qualitative differences or relationships between reference images, and their setups remain similar to those of conventional single-image editing benchmarks.
+Moreover, although designed for multi-reference generation, they do not account for the heterogeneous nature of reference images.
+This challenge is analogous to instruction following in LLMs [97, 35, 24, 23, 40], where performance degrades as the number of constraints grows; multi-reference image generation similarly requires models to satisfy all given references without neglecting any.
 Figure 2: Construction pipeline for our benchmark, consisting of four stages: (1) collecting high-quality real and synthetic images, (2) filtering out inappropriate or low-quality samples, (3) performing hierarchical category classification, and (4) generating and validating editing instructions by Gemini and humans.
@@
 In reference-driven image generation, tasks can be categorized based on the number and relational structure of reference images.
-In this work, we follow the overall design philosophy of prior benchmarks such as DreamOmni2 [xia2025dreamomni2], but place a stronger emphasis on an aspect that has been largely underexplored so far: the diversity and compositional relationships among multiple references.
-While existing benchmarks mainly evaluate performance in terms of the number of references or single-image editing accuracy, our benchmark explicitly aims to measure a model’s ability to understand and reason about inter-reference variations as a key part of multimodal generation.
+In this work, we follow the overall design philosophy of prior benchmarks such as DreamOmni2 [83], but extend the number of references up to 8 to enable a more fine-grained analysis of scalability.
+We also go beyond task-category-based evaluation by incorporating axes that capture the intrinsic properties of the reference set itself, such as domain gaps between references, scale mismatches, rare concepts, and multilingual text rendering.
 Single Reference.
-This is the most fundamental setting, corresponding to the standard image editing task.
+This is the simplest setting, corresponding to the standard image editing task.
 Given a single reference image and a textual instruction, the model generates an edited image that maintains semantic consistency and preserves the subject’s visual identity.
-This configuration follows traditional instruction-driven editing benchmarks such as MagicBrush [zhang2023magicbrush] and EMU-Edit [sheynin2024emuedit], which represent the foundation of reference-driven generation.
+This configuration follows traditional instruction-driven editing benchmarks such as MagicBrush [93] and EMU-Edit [69], which represent the foundation of reference-driven generation.
 Two References – 11 Task Types.
@@
 Multi References – Compositional Generation and Inter-Reference Relationships.
-The multi-reference setting uses 3–8 reference images, requiring the model to perform more compositional and reasoning-based generation.
-While prior works, such as DreamOmni2 [xia2025dreamomni2], mainly focus on compositional generation using multiple references, our benchmark takes a step further by making inter-reference diversity, consistency, and structural relations central to the evaluation.
-We define four base task structures, each evaluated with $\{3,4,5,6,7,8\}$ reference images, and include variants with and without textual instructions — resulting in a total of 48 tasks:
+The multi-reference setting uses 3–8 reference images, requiring the model to perform more compositional generation.
+We define four base task structures, each evaluated with $\{3,4,5,6,7,8\}$ reference images, resulting in 24 tasks:
   • X Objects: Compose $X$ objects from the reference set.
@@
   • X–1 Objects + Global: Arrange X1X-1 objects and modify the overall tone according to another reference.
-This design enables us to assess whether a model can effectively integrate multiple reference sources into a coherent, contextually consistent scene.
+This design enables us to assess whether a model can integrate multiple reference sources into a consistent scene.
 Diversity Factors in Multi-Reference Generation.
-A central contribution of our benchmark is the explicit incorporation of inter-reference diversity factors, which have been largely overlooked in previous benchmarks such as DreamOmni2.
-We highlight the following new evaluation dimensions:
-  • Number of references: systematic comparison across 3–8 inputs.
+A central contribution of our benchmark is the explicit incorporation of inter-reference diversity factors, which have been largely overlooked in previous benchmarks such as DreamOmni2 [83].
+We highlight the following dimensions:
   • Cross-domain diversity: mixing references from different domains such as real photos, anime, CG, or sketches.
@@
   • Multilingual references: assessing multilingual instruction understanding during rendering.
-These factors go beyond conventional image-editing evaluation and are critical for assessing a model’s compositional reasoning, identity preservation, and cross-domain generalization.
-In summary, our benchmark diverges from the conventional notion of multi-reference generation, such as in DreamOmni2, which primarily focuses on simple reference composition.
-Instead, it introduces a new framework designed to evaluate model capabilities to interpret relationships among heterogeneous references and integrate them into coherent generations.
+These factors go beyond prior benchmarks that focus on simple reference composition; our benchmark evaluates whether models can interpret relationships among heterogeneous references and integrate them into coherent outputs.
 Figure 3: (Left) Comparison between the statistics of real data only and those after adding synthetic data.
 The original dataset was biased toward background images, with few person- and object-related samples.
-To correct this imbalance, we generated additional synthetic images using Nanobanana and ChatGPT-Image-1, focusing on clear subjects such as people, animals, and objects.
-This significantly increased person- and object-related categories, resulting in a more balanced and comprehensive benchmark.
+To correct this imbalance, we generated additional synthetic images using Nano Banana and GPT-Image-1, focusing on clear subjects such as people, animals, and objects.
+This results in a more balanced and comprehensive distribution of data.
 (Right) Examples of synthesized images in each category.
 3.2 How to Construct MultiBanana
+The construction pipeline is illustrated in Figure 2.
 Image Collection.
 The reference images used in our benchmark consist of both real and synthetic data.
-All real images were extracted from the LAION-5B dataset [schuhmann2022laion].
-To ensure high quality, we selected only images with an Aesthetic score higher than 6.25 and a resolution greater than 512 pixels.
+All real images were extracted from the LAION-5B dataset [68].
+To ensure high quality, we selected only images with an Aesthetic score [43] higher than 6.25 and a resolution greater than 512 pixels.
 However, the collected real images exhibited a distributional bias (Figure 3; left).
-Specifically, landscape scenes appeared frequently, whereas categories useful as editing references, such as humans and objects, were underrepresented.
-While landscape images can serve as useful background references, the scarcity of human and object references limits the diversity of editing tasks that can be constructed in the benchmark.
-To address this imbalance, we generated additional synthetic images under diverse conditions using Nano Banana [google2025nanobanana] and ChatGPT-Image-1 [openai2025gpt4oimage], and used them as supplementary reference images (Figure 3; right).
-This approach not only mitigated the categorical bias in the reference set but also expanded the coverage and diversity of reference images by incorporating both real and synthetic examples.
+Specifically, landscape scenes appeared frequently, whereas humans and objects were underrepresented.
+While landscape images can serve as useful background references, the scarcity of human and object references limits the diversity of editing tasks.
+To address this imbalance, we generated additional synthetic images under diverse conditions using Nano Banana [22] and GPT-Image-1 [56], and used them as supplementary reference images (Figure 3; right).
+This approach not only mitigated the categorical bias in the reference set but also expanded its coverage and diversity by incorporating both real and synthetic examples.
 Image Filtering.
-Among the collected images, some contained small subjects that were difficult to use as references, or were inappropriate as reference images, such as those containing harmful content, tables, or corrupted data.
+Among the collected images, some contained small subjects that were difficult to use as references, or were inappropriate for reference use, such as images containing harmful content, tables, or corrupted data.
 To filter out such images and retain only those suitable for a reference-based image editing benchmark, we applied the following procedure.
-Following [ye2025imgedit], we extracted images that contained large and clearly visible objects suitable for editing tasks.
-Object detection was performed using YOLOv12 [tian2025yolov12] and SAM2 [ravi2025sam], while CLIP [radford2021clip] was used to verify semantic consistency.
+Following [91], we extracted images that contained large and clearly visible objects suitable for editing tasks.
+Object detection was performed using YOLOv12 [73] and SAM2 [64], while CLIP [62] was used to verify semantic consistency.
 Additionally, harmful or corrupted images were removed through a combination of automatic filtering with Gemini and manual inspection, resulting in the final curated set of real images.
-Figure 4: (Left) Breakdown of the multi-reference tasks.
-The editing sets were selected to ensure that the number of sets within each task is balanced across different reference counts.
-(Middle) For every X–references task, the dataset contains at least 390 editing sets.
-Further, each colored task category also includes at least 70 sets, which is larger than the prior work [xia2025dreamomni2].
-(Right) Word cloud generated from all prompts.
-It primarily consists of terms that describe a wide range of object categories as well as words indicating spatial directions.
 Category Classification.
-To construct multi-reference image generation tasks, we categorized each reference image accurately.
+To construct multi-reference image generation tasks, we categorized each reference image.
 For example, in a portrait transformation task that modifies a person’s hairstyle or makeup, both reference images must contain a person.
@@
 To satisfy such task-dependent conditions, we performed hierarchical image classification on a combined set of real and synthetic images.
-We first obtained automatic multi-level labels using Gemini and then conducted human verification to ensure correct categorization.
-At the top level, we defined five major categories: person, object, background, light, tone.
+We first obtained automated multi-level labels using Gemini, then conducted human verification to ensure correct categorization.
+At the top level, we defined six major categories: person, object, background, light, tone, and style.
 Each of these top-level categories was further divided into more specific subcategories.
 For instance, under the person category, we included male, female, hairstyle, makeup, pose useful, and face expression; and under the object category, we defined animal, clothing design, industrial design, and object.
-This hierarchical labeling enabled precise task construction and increased the benchmark’s diversity.
+This hierarchical labeling enabled systematic task construction.
 Task Construction.
 The editing instructions were generated using Gemini.
-For each task, we manually selected suitable reference categories, randomly sampled images, and presented them to Gemini to produce corresponding editing instructions.
+For each task, we manually selected suitable reference categories, randomly sampled images, and presented them to Gemini to generate corresponding editing instructions.
 Gemini then evaluated whether following each instruction would lead to visual breakdowns.
 For instance, in a replacement task, whether the replaced object would appear unnaturally suspended in mid-air.
-Among the generated candidates, we retained only the reference–instruction pairs that did not exhibit visual breakdowns and then asked Gemini to further assess whether each instruction constituted a challenging editing scenario, such as cross-domain modifications.
-Finally, we manually reviewed the remaining data to remove any residual cases that Gemini failed to detect or instructions that were deemed too trivial.
+Among the generated candidates, we retained only the reference–instruction pairs that did not exhibit visual breakdowns and then asked Gemini and human annotators to further filter out tasks that were nonsensical or too trivial.
 Difficult Reference Categorization.
@@
 The rare-concept category contained samples featuring animals with unusual colors or patterns, humans or animals with atypical makeup or clothing, or fictional creatures.
-The multilingual category included samples with text in two different languages among English, Chinese, and Japanese.
-Samples not matching any category were grouped as others.
-Table 2: Difficult reference combination statistics and ratios relative to the total 3769 tasks.
-Our benchmark provides a sufficient number of samples for assessing a model’s compositional reasoning, identity preservation, and cross-domain generalization, compared to prior work [xia2025dreamomni2].
+The multilingual category included samples with text in two different languages, including English, Chinese, and Japanese.
+Figure 4: (Left) Breakdown of the multi-reference tasks.
+The editing sets were selected to ensure that the number of sets per task is balanced across reference counts.
+(Middle) For every X–references task, the dataset contains at least 390 editing sets.
+Furthermore, each colored task category includes at least 70 sets, which is larger than in prior work [83].
+(Right) Word cloud generated from all prompts.
+It primarily consists of terms that describe a wide range of object categories and words indicating spatial directions.
+3.3 Statistics in MultiBanana
+Image Category Statistics.
+This section presents a statistical overview of both the real and synthetic source data used to construct our benchmark.
+The dataset is organized into six major categories and thirteen subcategories based on visual and semantic attributes.
+In the real dataset, the largest major category is background, comprising 798 images, one-fourth of the total.
+These mainly consist of images without distinct subjects.
+The next-largest categories are person (741 images) and style (579 images), and together these three categories account for 70% of the dataset.
+In contrast, light (195 images), object (413 images), and tone (118 images) are small in number.
+Although these categories are useful as references for changing global conditions such as illumination and overall atmosphere, their sample sizes remain limited.
+Within the object category, there are 111 images of animals, 40 of clothing and textile design, 71 of industrial design, and 191 of miscellaneous objects, indicating that data focusing on specific materials or design patterns is scarce.
+Most images in the background, style, and tone categories lack clear subjects.
+Furthermore, images containing textual elements, such as logos or signs, are nearly absent.
+The person category is subdivided into finer attributes: face expression (10), female (181), male (139), hairstyle (135), makeup (59), and pose usefulness (217).
+While pose-related data account for the largest portion (around 30%), reference data capturing subtle characteristics such as facial expression or makeup remain limited, resulting in a skewed distribution.
+To address this imbalance, synthetic data were generated, focusing primarily on images containing explicit subjects (e.g., persons, animals, and objects) and on those including text.
+The generation process was designed to intentionally expand the coverage of underrepresented subcategories and improve the overall balance among attributes.
+[…]
+Consequently, the dataset became more balanced between background-centric and subject-centric images, with broader coverage of diverse visual attributes.
+This improvement enhances the benchmark’s comprehensiveness and enables more versatile evaluation and instruction-based generation tasks under a wide range of conditions.
+Task Statistics.
+After filtering the editing instructions using both Gemini and manual checks, we obtained 3,769 high-quality reference images and instruction sets.
+Among them, 264 sets correspond to single-reference tasks, 907 to two-reference tasks, and 2,598 to multi-reference tasks.
+Figure 4 provides a further breakdown of the multi-reference tasks, showing that they are evenly distributed with respect to the number of reference images.
+Additionally, for each task type, X Object, X–1 Object + Background, X–1 Object + Local, and X–1 Object + Global, we ensured that at least 70 sets were included.
+This is larger than the number of sets in multi-reference tasks of OmniContext [80] and DreamOmni2 [83], enabling more reliable benchmarking.
+Furthermore, Table 2 shows that our newly introduced evaluation axes – cross-domain, scale and viewpoint differences, rare concept, and multilingual – provide sufficient samples.
+See Appendix C for details.
+Table 2: Statistics and ratios of difficult reference combinations relative to the total 3,769 tasks.
+Our benchmark provides sufficient samples for assessing a model’s capacity to interpret relationships among heterogeneous references and integrate them into outputs.
 Cross Different Rare Multi-
@@
 Table 3: Average performance per model across different task categories (10-point scale; the higher the better).
-The average scores from Gemini-2.5 and GPT-5 are reported.
+The average scores from Gemini 2.5 and GPT-5 are reported.
 Both open-source and closed-source models exhibit notably lower performance on background replacement tasks, regardless of the number of reference images.
@@
 OmniGen2 5.919 3.442 3.256 3.598 3.369 3.022
-3.3 Statistics in MultiBanana
-Image Category Statistics.
-This section presents a statistical overview of both the real and synthetic source data used to construct our benchmark.
-The dataset is organized into six major categories and thirteen subcategories based on visual and semantic attributes.
-In the real dataset, the largest major category is background, comprising 798 images, one-fourth of the total.
-These mainly consist of images without distinct subjects.
-The next largest categories are person (741 images) and style (579 images), and together these three categories account for 70% of the dataset.
-In contrast, light (195 images), object (413 images), and tone (118 images) are relatively small in size.
-Although these categories are useful as references for changing global conditions such as illumination and overall atmosphere, their sample sizes remain limited.
-Within the object category, there are 111 images of animals, 40 of clothing and textile design, 71 of industrial design, and 191 of miscellaneous objects, indicating that data focusing on specific materials or design patterns is scarce.
-Most images in the background, style, and tone categories lack clear subjects.
-Furthermore, images containing textual elements, such as logos or signs, are nearly absent.
-The person category is subdivided into finer attributes: face expression (10), female (181), male (139), hairstyle (135), makeup (59), and pose usefulness (217).
-While pose-related data account for the largest portion (around 30%), reference data capturing subtle characteristics such as facial expression or makeup remain limited, resulting in a skewed distribution.
-To address this imbalance, synthetic data were generated, focusing primarily on images containing explicit subjects (e.g., persons, animals, and objects) as well as those including text.
-The generation process was designed to intentionally expand the coverage of underrepresented subcategories and improve the overall balance among attributes.
-[…]
-Consequently, the dataset became more balanced between background-centric and subject-centric images, with broader coverage of diverse visual attributes.
-This improvement enhances the benchmark’s comprehensiveness and enables more versatile evaluation and instruction-based generation tasks under a wide range of conditions.
 Figure 5: Changes in scores for each evaluation criterion when varying the number of reference images.
-Both open-source and closed-source models exhibit a general trend of decreasing all scores as the number of references increases.
-Task Statistics.
-After filtering the generated editing instructions using both Gemini and manual checks, we obtained 3,769 high-quality reference images and instruction sets.
-Among them, 264 sets correspond to single-reference tasks, 907 to two-reference tasks, and 2,598 to multi-reference tasks.
-Figure 4 provides a further breakdown of the multi-reference tasks, showing that they are evenly distributed with respect to the number of reference images.
-Additionally, for each task type, X Object, X–1 Object + Background, X–1 Object + Local, and X–1 Object + Global, we ensured that at least 70 sets were included.
-This is substantially larger than the number of sets in multi-reference tasks of DreamOmni2, which is on the order of ten, enabling more reliable evaluation of model performance.
-Furthermore, Table 2 shows that our newly introduced evaluation axes – cross-domain, scale and viewpoint differences, rare concept, and multilingual – provide sufficient samples.
-See Appendix B for details.
-Figure 6: Results for diffucult reference combinations. For cross-domain and different scale and view tasks, every model shows lower scores than tasks without such conditions.
+Both open-source and closed-source models exhibit a general trend of decreasing scores across all metrics as the number of references increases.
 3.4 Evaluation Setting
 For evaluating the quality of edited images, human evaluation is among the most valuable methods, yet gathering human feedback at scale is prohibitively costly.
-A practical approach to reduce time and cost is to leverage AI feedback from VLMs (e.g., Gemini-2.5 [geminiteam2023gemini] and GPT-5 [openai2025gpt5]), which has been shown to modestly align with human judgments of image quality [na2024boost, wu2024boosting, furuta2024improving, oshima2025inference].
-Therefore, we used Gemini-2.5 and GPT-5 to rate all generated images.
+A practical approach to reducing time and cost is to leverage AI feedback from VLMs (e.g., Gemini 2.5 [19] and GPT-5 [57]), which has been shown to align reasonably with human judgments of image quality [38, 53, 82, 17, 58].
+Therefore, we used Gemini 2.5 and GPT-5 to rate all generated images.
 […]
 4 Experiments
-We adopt several closed models (i.e., Nano Banana [google2025nanobanana], GPT-Image-1 [openai2025gpt4oimage]) and open models (i.e., Qwen-Image-Edit-2509 [wu2025qwenimagetechnicalreport]], DreamOmni2 [xia2025dreamomni2], OmniGen2 [wu2025omnigen2]) for MultiBanana evaluation.
+We adopt several closed models (i.e., Nano Banana [22], GPT-Image-1 [56]) and open models (i.e., Qwen-Image-Edit-2509 [79], DreamOmni2 [83], OmniGen2 [80]) for MultiBanana evaluation.
 For all models except OmniGen2, we generate outputs for every combination of reference images and text prompts included in the MultiBanana benchmark.
-Since OmniGen2 does not support more than six reference images, we only evaluate it under the 5-reference setting.
-As described in Section 3.4, the evaluation is conducted using Gemini-2.5 [geminiteam2023gemini] and GPT-5 [openai2025gpt5].
-In the main paper, we report the average scores from Gemini-2.5 and GPT-5, while the individual results for each evaluator are provided in the supplementary material.
-Figure 7: (Left) Failure cases of cross-domain, different scale, rare concept, multilingual text.
-(Right) Failure cases of multi-references in open-source models.
-They tend to ignore multiple subjects under high-reference conditions.
+Since OmniGen2 does not support more than 6 reference images, we evaluate it only under the 5-reference setting.
+As described in Section 3.4, Gemini 2.5 [19] and GPT-5 [57] are used for evaluation.
+In the main paper, we report the average scores from Gemini 2.5 and GPT-5.
+See Appendix A for the results of Qwen3-VL as judge, and Appendix F for the detailed results and discussion.
+Figure 6: (Left) Failure cases of closed-source models under challenging conditions such as cross-domain references, scale mismatches, rare concepts, and multilingual text.
+(Right) Failure cases of open-source models under many-reference conditions.
+They tend to ignore some reference subjects entirely.
 4.1 Per-Task Evaluation
-We compute the MultiBanana scores for each task type defined in Section 3.1, and present the results in Table 3 and Section E.1.
+We compute the MultiBanana scores for each task type defined in Section 3.1, and present the results in Table 3.
 For the two-reference tasks, we report the average score across all 11 task variants.
 For the multi-reference tasks, we compute the average score for each task type (i.e., X Objects, X–1 Objects + Local, X–1 Objects + Global, X–1 Objects + Background) across $X=\{3,4,5,6,7,8\}$.
-Overall, both open-source and closed-source models exhibit notably lower performance on background replacement tasks, regardless of the number of reference images.
-The results for the two-reference setting also reveal clear performance gaps between closed-source and open-source models, even for relatively simple editing scenarios such as single-reference or two-reference tasks.
+Both open-source and closed-source models exhibit lower performance on background-replacement tasks, regardless of the number of reference images.
+The results also reveal clear performance gaps between closed-source and open-source models, even in relatively simple editing scenarios, such as single- and two-reference tasks.
 4.2 Effect of the Number of References
 We investigate how the MultiBanana score changes when varying the number of reference images $X=\{3,4,5,6,7,8\}$ for each multi-reference generation task (Figure 5).
-Both open-source and closed-source models exhibit a general trend of decreasing Total Score as the number of references increases.
-However, the nature of this degradation differs markedly between the two model families.
+Both open-source and closed-source models exhibit a general trend of decreasing Total Score with increasing number of references.
+However, the nature of this degradation differs between the two model families.
 Closed-source models demonstrate strong adherence to reference attributes and prompt instructions.
 As a result, even when the number of references grows, the decline in Instruction Alignment and Reference Consistency remains relatively mild.
-Nevertheless, because these models attempt to satisfy all reference constraints strictly, increasing the number of subjects often leads to degraded overall visual quality, including over-crowded scenes and compositional inconsistencies (see Figure 1 and Figure 7; left).
-In contrast, open-source models show weaker adherence to prompts and references; even under the 3-reference setting, both Instruction Alignment and Reference Consistency remain low, and these scores further deteriorate as the reference count increases.
+Nevertheless, because these models attempt to strictly satisfy all reference constraints, increasing the number of subjects often degrades overall visual quality, including overcrowded scenes and compositional inconsistencies (see Figure 1 and Figure 6; left).
+In contrast, open-source models exhibit weaker adherence to prompts and references; even under the 3-reference setting, both Instruction Alignment and Reference Consistency remain low, and these scores deteriorate further as the number of references increases.
 Qualitative inspection also reveals that open models frequently ignore multiple subjects under high-reference conditions.
 However, their overall quality (background consistency, physical realism, visual quality) remains relatively stable, and they tend not to suffer from the severe compositional collapse observed in closed models (see Figure 6; right).
-In summary, closed-source models succeed in incorporating all required subjects but often produce globally inconsistent or distorted scenes due to overfitting to fine-grained reference details.
+In summary, closed-source models succeed in incorporating all required subjects but often generate globally inconsistent or distorted scenes due to overfitting to fine-grained reference details.
 Open-source models, by contrast, tend to omit multiple reference subjects in complex scenarios but generate visually clean images with fewer structural failures.
 These observations highlight a clear trade-off between strict reference fidelity and holistic image coherence.
+Figure 7: Results for difficult reference combinations. For cross-domain and different scale and view tasks, every model shows lower scores than tasks without such conditions.
 4.3 Analysis of Difficult Reference Combinations
-Next, to analyze the challenging subject combinations included in this benchmark, namely those described in Section 3.1, we compare the MultiBanana scores between tasks containing these difficult combinations and those that do not in Figure 6.
-For cross-domain, different-scale and view, rare-concept, and multilingual tasks, most models show lower scores than for tasks without such conditions.
+To analyze the challenging reference combinations described in Section 3.1, we compare MultiBanana scores between tasks containing these difficult combinations and those that do not, as shown in Figure 7.
+For cross-domain, different scale and view, rare-concept, and multilingual tasks, most models achieve lower scores than on tasks without such conditions.
 These results indicate that the difficult reference concepts focused on by our benchmark remain challenging, even for state-of-the-art closed models.
 This demonstrates that our benchmark effectively uncovers the current limitations of multi-reference image generation.
+4.4 Reliability of VLM Judge
+To assess the reliability of VLM-based evaluation, we measured the correlation between VLM and human ratings using three human evaluators on 32 samples from our benchmark (Table 4).
+Both GPT-5 and Gemini 2.5 show strong correlation with human judgments.
+This validates the use of VLM judges as a time- and cost-efficient proxy for human evaluation, as described in Section 3.4.
+We also confirmed that Qwen3-VL-8B-Instruct [2] is a reliable open-source alternative, thereby broadening accessibility for researchers without API access.
+See Appendix E for further discussion.
 4.5 Agentic Framework
@@
 We then conduct experiments on our MultiBanana benchmark.
-We show the results of IPR in Table 3, where the generator creates images, and based on these images, the planner rewrites the prompt over multiple steps.
+We show the results of IPR in Table 3, where the generator creates images, and the planner rewrites the prompt over multiple steps based on these images.
 While GPT’s performance improved with the agent framework, Nano Banana showed no improvement and, in some cases, performance degraded due to information loss from the initial instruction.
-For details on the problem setting, ablations, and results, see the supplementary material.
+For details on the problem setting, ablations, and results, see Appendix F.7.
+Table 4: Correlation between human and VLM judges.
+Judge Pearson $r$ Spearman $r$ Cohen’s $\kappa$
+GPT-5 0.6936 0.7134 0.6135
+Gemini 2.5 0.5726 0.6113 0.3908
+Qwen3-VL 0.5457 0.5642 0.5235
 5 Conclusion
 We constructed a comprehensive benchmark for multi-reference image generation, focusing on challenges specific to this setting, such as difficult reference combinations.
-Using it, we evaluated state-of-the-art open and closed models and found two notable failure modes as the number of references increases: models either ignore editing instructions or collapse when trying to follow them.
+Using it, we evaluated state-of-the-art open and closed models and found two notable failure modes as the number of references increased: models either ignored editing instructions or collapsed when trying to follow them.
 Across different types of reference difficulty, tasks involving cross-domain inputs, large-scale or viewpoint gaps, rare concepts, and multilingual inputs are consistently challenging.
 Acknowledgements
-We thank Minsu Kim and Heiga Zen for their support on this work and review on the initial version of the paper.
-We also appriciate the funding support from Google Japan.
+We thank Minsu Kim and Heiga Zen for their support with this work and for their review of the initial version of the paper.
+We also appreciate the funding support from Google Japan.
 MS was supported by JSPS KAKENHI Grant Number JP23H04974.
-[…]
 Supplementary Material
-A Details on Benchmark Construction
-A.1 Details on Image Collection
-We constructed the reference image collection for our benchmark using both real images sourced from LAION-5B [laion2022aesthentic] and synthetic images generated by Nano Banana [google2025nanobanana] and GPT-Image-1 [openai2025gpt4oimage].
-By combining real and synthetic data, we mitigated the categorical biases inherent in the real-image distribution while substantially expanding the coverage and diversity of the reference set.
+A Results of Qwen3-VL Judge
+Table 5: Per-task total scores for each single- and two-reference task, evaluated by Qwen3-VL.
+“Back.” denotes the Background task.
+Both open-source and closed-source models exhibit lower performance on background replacement tasks.
+Note that for Nano Banana and GPT-Image-1, we use the versions available as of January 2026.
+Model Single Add Back.
+Color Hair Makeup Material Pose Replace Style Text Tone DreamOmni2 8.47 6.51 4.59 8.60 7.40 7.89 6.76 6.25 6.86 5.10 5.48 5.14 OmniGen2 9.04 5.34 6.04 5.07 6.70 6.68 6.12 5.91 6.34 3.24 4.29 4.00 Qwen-Image-Edit-2509 9.16 7.01 4.34 6.46 8.00 6.68 5.86 5.87 7.57 3.41 4.95 6.08 Nano Banana 9.45 8.13 7.97 9.67 9.16 9.25 8.95 8.76 8.30 9.03 8.81 8.52 GPT-image-1 9.51 8.09 8.08 9.70 9.04 9.24 9.17 8.80 8.30 9.10 8.52 8.78
+Table 6: Per-task total scores for each multi-reference task, evaluated by Qwen3-VL.
+The local, global, back, and object columns under X correspond to X–1 Objects + Local, X–1 Objects + Global, X–1 Objects + Background, and X Object, respectively.
+Both open-source and closed-source models exhibit a general trend of decreasing scores across all tasks as the number of references increases.
+They also tend to perform worse on background replacement tasks, especially as the number of reference images increases.
+Note that for Nano Banana and GPT-Image-1, we use the versions available as of January 2026.
+Model 3-references 4-references 5-references local global back object local global back object local global back object DreamOmni2 4.39 4.27 4.33 3.68 3.80 2.95 2.77 2.90 3.23 2.12 2.86 2.26 OmniGen2 4.71 4.86 4.39 4.13 4.37 3.86 3.43 2.95 3.99 2.75 2.14 2.26 Qwen-Image-Edit-2509 5.22 5.81 5.81 5.63 3.24 4.51 4.42 3.94 1.05 1.48 1.28 1.10 Nano Banana 7.63 7.21 6.96 6.91 7.96 6.77 6.47 6.78 7.96 6.54 6.03 5.87 GPT-image-1 7.83 7.26 6.99 6.72 7.73 6.74 5.94 6.83 7.92 6.54 6.16 5.83 Model 6-references 7-references 8-references local global back object local global back object local global back object DreamOmni2 3.34 2.26 1.72 1.93 2.61 2.45 2.10 1.81 2.44 1.65 1.83 1.59 OmniGen2 - - - - - - - - - - - - Qwen-Image-Edit-2509 1.00 1.49 1.00 1.00 1.00 1.46 1.00 1.13 1.00 1.00 1.00 1.00 Nano Banana 6.84 6.08 5.34 5.64 6.37 5.78 5.13 5.44 6.46 6.24 5.31 5.28 GPT-image-1 7.11 5.86 5.51 5.92 6.81 5.78 5.37 5.74 7.07 5.86 5.14 5.55
+In Section 4, we report the average scores from Gemini 2.5 and GPT-5 as judges.
+As demonstrated in Section 4.4, Qwen3-VL-8B-Instruct [2] achieves strong correlation with human judgments comparable to GPT-5 [57] and Gemini 2.5 [19], confirming its reliability as a judge model (see Table 4).
+Although we used stable API versions for the closed-source judges, to ensure full reproducibility, we additionally employ Qwen3-VL-8B-Instruct, an open-weight, fixed-version model, as the judge.
+This also broadens access to our benchmark for researchers who do not have access to closed-source VLMs via APIs.
+Table 5 and Table 6 present per-task total scores evaluated using Qwen3-VL as the sole judge, calculated in the same way as described in Section 3.4.
+The results are broadly consistent with those obtained using closed-model judges.
+The overall ranking of image generation models is preserved.
+We also observe the same key trends: performance degrades as the number of reference images increases, and background replacement tasks tend to yield lower scores across both open-source and closed-source models.
+These observations reinforce the validity of Qwen3-VL as a reliable, reproducible alternative.
+We expect this fixed, open-weight judge to serve as a dependable baseline for future evaluations on our benchmark.
+B Details on Benchmark Construction
+B.1 Image Collection
+We constructed the reference image collection for our benchmark using both real images sourced from LAION-5B [68] and synthetic images generated by Nano Banana [22] and GPT-Image-1 [56].
+By combining real and synthetic data, we mitigated the categorical biases inherent in the real-image distribution while expanding the coverage and diversity of the reference set.
 To generate synthetic images, we designed distinct axes of variation for four major categories: humans, animals, objects, and text-containing images.
-Human Category  We systematically combined attributes such as makeup, hairstyle, hair color, pose, strong facial expressions, lighting, and overall tonal style to create a wide range of appearance conditions.
-Examples include ethnic or fashion-model-style makeup, vivid or unconventional hair colors, model-like or manga-like poses, strong emotional expressions (e.g., anger, joy, surprise), dramatic lighting such as window light through blinds or neon illumination, and global color tones such as sepia or bluish palettes.
-These conditions were intentionally selected to supplement variations that are underrepresented in real-world data.
-Animal Category  We first specified major species groups (dogs, cats, birds, reptiles, wild animals, and fantasy animals).
-Then we added attributes to enhance the diversity specific to animal imagery.
+Human Category.
+We systematically combined attributes such as makeup, hairstyle, hair color, pose, strong facial expressions, lighting, and overall tonal style to create a wide range of appearance conditions.
+Examples include ethnic or fashion-model-style makeup, vivid or unconventional hair colors, model-like or manga-like poses, strong emotional expressions (e.g., anger, joy, surprise), dramatic lighting (e.g., window light through blinds or neon illumination), and global color palettes (e.g., sepia or bluish).
+These conditions were intentionally selected to supplement the variations underrepresented in real-world data.
+Animal Category.
+We first specified major species groups (dogs, cats, birds, reptiles, wild animals, and fantasy animals).
+We then added attributes to enhance the diversity of animal imagery.
 These included human-like behaviors (e.g., a bear reading a book, a fox playing in a rock band), rare colors (e.g., a blue horse or a pink panda), distinctive lighting conditions, and varied global tones such as sepia.
 This incorporated rare or extreme examples that seldom appear in real datasets.
-Object Category  We expanded diversity along two axes: intrinsic object attributes and photographic conditions.
+Object Category.
+We expanded diversity along two axes: intrinsic object attributes and photographic conditions.
 The former included distinctive textures (e.g., denim, leather, knit), characteristic prints (e.g., floral, striped, checkered), and variations in material or design (e.g., matte, glossy, transparent).
 The latter included variations in lighting and global tonal adjustments (e.g., sepia, cool color tones), thereby ensuring a broad range of appearance conditions for objects.
-Images Containing Text  Images containing text, which were largely absent from real data, were generated exclusively using GPT-Image-1, as Nano Banana exhibited limited ability to render non-English text.
+Images Containing Text.
+Images containing text, which were largely absent from real data, were generated exclusively using GPT-Image-1, as Nano Banana exhibited limited ability to render non-English text.
 We defined three levels of text length (a single word, a short phrase of up to five words, and a longer sentence).
 We combined each with two scene types: text rendered directly within a landscape (e.g., ocean, forest) and text appearing naturally in the environment (e.g., on signs in real scenes).
-These six prompt patterns were generated in English and subsequently translated into Japanese and Chinese to enrich multilingual text-containing images.
-Scene Diversity  Finally, all synthetic prompts were designed to incorporate scene diversity, including European cityscapes, university campuses, tropical beaches, event halls, and other varied environments.
-The generated subjects were ensured to be well-framed, resulting in images suitable for use as high-quality reference material.
-A.2 Details on Image Filtering
-First, to remove images unsuitable for editing, the images were filtered based on the size of the foreground objects, following the manner in [ye2025imgedit].
-YOLOv12 [tian2025yolov12] was applied to all images to detect objects and then performed semantic segmentation using SAM [ravi2025sam], conditioned on the detected bounding boxes.
-Images were discarded if they met any of the following criteria: the bounding box area covered less than 2% of the entire image, the segmented region within the bounding box covered less than 30%, or the CLIP similarity [ma2022xclip] between the YOLO-predicted class name and the bounding box region was below 20.
+These six prompt patterns were generated in English and subsequently translated into Japanese and Chinese to enrich the multilingual text in images.
+Scene Diversity.
+Finally, all synthetic prompts were designed to incorporate scene diversity, including European cityscapes, university campuses, tropical beaches, event halls, and other varied environments.
+The generated subjects were well-framed, resulting in images suitable for use as high-quality reference material.
+Figure 8: Examples of a duplicated synthetic image.
+B.2 Image Filtering
+First, to remove images unsuitable for editing, we filtered them based on the size of the foreground objects, as in [91].
+YOLOv12 [73] was applied to all images to detect objects and then performed semantic segmentation using SAM [64], conditioned on the detected bounding boxes.
+Images were discarded if they met any of the following criteria: the bounding box area covered less than 2% of the entire image, the segmented region within the bounding box covered less than 30%, or the CLIP similarity [48] between the YOLO-predicted class name and the bounding box region was below 20.
 These cases were judged to contain objects that were not clearly visible.
@@
 After this filtering, 48% of the images were retained.
-We also filtered out images that were inappropriate or unsuitable as references, such as unsafe content, charts, and screenshots of system messages, using both Gemini-based automated screening and human review.
+We also filtered out images that were inappropriate or unsuitable as references, such as unsafe content, charts, and screenshots of system messages, using both automated screening based on Gemini and human review.
 For inappropriate content, we specifically targeted categories including hate, harassment, violence, self-harm, sexual content, nudity, shocking content, illegal activity, and other distressing material, ultimately excluding approximately 3% of the collected images.
-We also manually removed synthetic data that was nearly identical.
+We also manually removed nearly identical synthetic data.
 An example is shown in Figure 8.
-Figure 8: Examples of a duplicated synthetic image.
 Figure 9: (Left) Number of tasks by reference count.
@@
 It contains eleven tasks.
-A.3 Details on Category Classification
+Figure 10: Example of “Hard” category of ImgEdit. These tasks are almost fully solvable by advanced models such as Nano Banana.
+Figure 11: Example of DreamOmni2 benchmark. These tasks are almost fully solvable by advanced models such as Nano Banana.
+B.3 Category Classification
 To construct multi-reference image editing tasks, each reference image had to be accurately categorized.
 For example, in portrait transformation tasks that modify a person’s hairstyle or makeup, both reference images must contain a person.
-Similarly, in replacement tasks involving humans or objects, the corresponding entities must be present in the reference images, and for background replacement tasks, background scenes must be determined in the reference images.
+Similarly, in replacement tasks involving humans or objects, the corresponding entities must be present in the reference images, and in background replacement tasks, the background scenes must be identified in the reference images.
 To satisfy these task-dependent requirements, we performed hierarchical image classification on a large combined set of real and synthetic images.
-For synthetic data, attributes can be specified directly during generation, eliminating the need for additional classification.
+For synthetic data, attributes can be specified directly in the prompt during generation, eliminating the need for additional classification.
 In contrast, real images exhibit substantial variability, requiring precise categorization to determine whether they can serve as valid reference images.
@@
 If there is precisely one person, the model additionally classifies attributes such as gender, makeup, hairstyle distinctiveness, pose clarity, and facial expression characteristics.
-The output follows a fixed template format, allowing consistent extraction of appearance features relevant to multi-reference generation.
+The output follows a fixed template, enabling consistent extraction of appearance features relevant to multi-reference generation.
 Object Image Classification  Object image classification follows a similar structure.
@@
 When exactly one object is detected, it is further categorized according to its visual and material properties, distinguishing among animal-like forms, patterned or decorative textiles, engineered or manufactured materials, and objects lacking distinctive design features.
-Background Image Classification  Background classification decomposes the scene into realism, tonal uniformity, and lighting strength.
+Background Image Classification  Background classification is conducted from three viewpoints: realism, tonal uniformity, and lighting strength.
 Realism distinguishes photographic scenes from stylized or synthetic ones; tone indicates whether the overall color distribution is uniform or varied; and lighting strength identifies whether a clear directional key light is present.
-Ambiguous lighting is conservatively labeled as absent to maintain annotation consistency.
-A.4 Details on Task Construction
+B.4 Task Construction
 The editing instructions were generated using Gemini.
@@
 Afterward, several instruction examples were provided to Gemini, and we asked Gemini to generate new instructions following the same syntax.
-After that, editing instructions that would cause visual breakdowns were removed by scoring them with Gemini.
-Additionally, the difficulty of editing instructions was also assessed by Gemini, and those judged easy, such as cases without domain differences, were removed.
+Afterward, editing instructions that could cause visual breakdowns were removed after being scored with Gemini.
+Additionally, Gemini assessed the difficulty of editing instructions, and those judged easy, such as cases without domain differences, were removed.
 Finally, the editing instructions were manually verified not to cause image breakdowns or result in overly simple edits, and assigned the remaining samples to the appropriate difficulty categories (cross-domain, scale and viewpoint differences, rare concept, and multilingual text rendering).
-B Further Statistics for MultiBanana
+C Further Statistics for MultiBanana
 The left of Figure 11 shows the number of tasks by each reference count.
@@
 The right of Figure 11 shows the breakdown of the two-reference tasks.
-It contains eleven tasks considered in the prior work [xia2025dreamomni2]: subject addition, subject replacement, background change, color modification, material modification, pose modification, hairstyle modification, makeup modification, tone transformation, style transfer, and text correction.
-Each task is ensured to include at least 6% of the editing samples, corresponding to roughly 60 samples, which is larger than the number of samples in xia2025dreamomni2.
-Figure 10: Qualitative example of “Hard” category of ImgEdit. These tasks are close to being fully solvable by advanced models such as Nano Banana.
-Figure 11: Qualitative example of DreamOmni2 benchmark. These tasks are close to being fully solvable by advanced models such as Nano Banana.
-C Further Comparison with Prior Benchmarks
-Table 1 shows that existing benchmarks do not provide systematic evaluation across diverse multi-reference conditions, support only a limited number of references, and fail to adequately account for important factors such as differences and compatibility among reference images.
-To illustrate that existing benchmarks are almost solved by state-of-the-art models such as Nano Banana [google2025nanobanana] and GPT-Image-1 [openai2025gpt4oimage], we evaluate the image generation capabilities of these models using ImgEdit [ye2025imgedit], the latest benchmark for image editing, and DreamOmni2 [xia2025dreamomni2], the state-of-the-art benchmark for multi-reference image generation.
-The quantitative results show that recent closed-source, state-of-the-art models achieve consistently high scores, indicating that these benchmarks are nearing saturation and may soon be unable to meaningfully distinguish among high-performing models (Table 4 and Table 5).
-The evaluation results are taken from the official GitHub repository of ImgEdit [^1] https://github.com/PKU-YuanGroup/ImgEdit?tab=readme-ov-file and the results reported in the DreamOmni2 paper.
-Our qualitative results suggest that, even in the additional “Hard” category of ImgEdit, most tasks are already close to being fully solvable by advanced models such as Nano Banana (Figure 10).
-A similar tendency is observed for DreamOmni2 benchmark as well (Figure 11).
-Taken together, these findings indicate that current benchmarks for image editing and multi-reference image generation are approaching their limits in evaluating cutting-edge models.
-There is a clear need for more challenging benchmarks that can more effectively capture the capabilities and failure modes of the latest generation of image models.
-Table 4: Comparison results of different models on ImgEdit-Bench [ye2025imgedit].
+It contains eleven tasks considered in the prior work [83]: subject addition, subject replacement, background change, color modification, material modification, pose modification, hairstyle modification, makeup modification, tone transformation, style transfer, and text correction.
+Each task is guaranteed to include at least 6% of the editing samples, corresponding to roughly 60 samples, which exceeds the number of samples in Xia et al.
+[83].
+Table 7: Comparison results of different models on ImgEdit-Bench [91].
 “Overall” is computed by averaging the scores across all task types.
-GPT-4.1 [openai2023gpt4] is used for evaluation.
+GPT-4.1 [54] is used for evaluation.
 The results show that recent closed-source, state-of-the-art models achieve substantially higher scores, suggesting that the benchmark may be approaching its ceiling in distinguishing high-end models.
-Evaluation results excluding Nano Banana are taken from the official GitHub repository of ye2025imgedit.
+Evaluation results excluding Nano Banana are taken from the official GitHub repository of Ye et al.
+[91].
 Model Add Adjust Extract Replace Remove Background Style Hybrid Overall
-MagicBrush [zhang2023magicbrush] 2.84 1.58 1.51 1.97 1.58 1.75 2.38 1.62 1.83
-AnyEdit [yu2025anyedit] 3.18 2.95 1.88 2.47 2.23 2.24 2.85 1.56 2.45
-OmniGen2 [wu2025omnigen2] 3.57 3.06 1.77 3.74 3.20 3.57 4.81 2.52 3.44
-Kontext-dev [labs2025flux1kontext] 3.83 3.65 2.27 4.45 3.17 3.98 4.55 3.35 3.71
-Nano Banana [google2025nanobanana] 3.63 4.66 3.73 4.69 4.71 4.60 4.47 4.10 4.37
-GPT-Image-1 [openai2025gpt4oimage] 4.61 4.33 2.90 4.35 3.66 4.57 4.93 3.96 4.20
-Table 5: Quantitative comparison of multimodal instruction-based editing and generation in DreamOmni2 Benchmark [xia2025dreamomni2].
-Both tasks are evaluated using Gemini [geminiteam2023gemini] and Doubao [bytedance2025doubao].
-This result shows that recent closed-source, state-of-the-art models achieve substantially higher scores, indicating that the benchmark may be reaching its ceiling for distinguishing among high-end models.
-We refer to evaluation results from xia2025dreamomni2.
+MagicBrush [93] 2.84 1.58 1.51 1.97 1.58 1.75 2.38 1.62 1.83
+AnyEdit [92] 3.18 2.95 1.88 2.47 2.23 2.24 2.85 1.56 2.45
+OmniGen2 [80] 3.57 3.06 1.77 3.74 3.20 3.57 4.81 2.52 3.44
+Kontext-dev [41] 3.83 3.65 2.27 4.45 3.17 3.98 4.55 3.35 3.71
+Nano Banana [22] 3.63 4.66 3.73 4.69 4.71 4.60 4.47 4.10 4.37
+GPT-Image-1 [56] 4.61 4.33 2.90 4.35 3.66 4.57 4.93 3.96 4.20
+Table 8: Quantitative comparison of multimodal instruction-based editing and generation in DreamOmni2 Benchmark [83].
+Both tasks are evaluated using Gemini [19] and Doubao [5].
+This result shows that recent closed-source, state-of-the-art models achieve substantially higher scores, suggesting that the benchmark may be approaching its ceiling in distinguishing among high-end models.
+We refer to evaluation results from Xia et al.
+[83].
 Method Editing Task Generation Task
@@
 Gemini Doubao Gemini Doubao Gemini Doubao Gemini Doubao
-Omnigen2 [wu2025omnigen2] 0.2195 0.2927 0.0427 0.0793 0.2083 0.2500 0.1000 0.0778
-Kontext [labs2025flux1kontext] 0.0488 0.1220 0.0183 0.0122 0.2500 0.3750 0.0556 0.1222
-Qwen-Image-Edit-2509 [wu2025qwenimagetechnicalreport] 0.2683 0.2927 0.0488 0.1159 0.1250 0.2917 0.1111 0.1556
-DreamOmni2 [xia2025dreamomni2] 0.5854 0.6585 0.5854 0.6280 0.5833 0.6667 0.5778 0.6333
-Nano Banana [google2025nanobanana] 0.6829 0.7073 0.6463 0.5488 0.5000 0.5417 0.5556 0.5488
-GPT-Image-1 [openai2025gpt4oimage] 0.6829 0.7805 0.7195 0.7439 0.6250 0.6250 0.6889 0.6333
-D Further Discussions for AI and Human Evaluation
-The correlation between AI and human evaluation has been shown in prior works [na2024boost, wu2024boosting, furuta2024improving, oshima2025inference].
-Consequently, using VLMs is becoming a standard method for the automatic evaluation of generated images [ye2025imgedit, wu2025omnigen2, xia2025dreamomni2].
-We used VLMs to evaluate the images along five metrics: Instruction Alignment, Reference Consistency, Background Subject Match, Physical Realism, and Visual Quality.
-The overall score is calculated as a weighted average of these metrics.
-The first two metrics represent the fundamental components of reference-based generation, while the last three metrics break down the overall quality.
-Therefore, the overall score is calculated as $\mathrm{avg}(\mathrm{Instruction\,Alignment}+\mathrm{Reference\,Consistency}+\mathrm{avg}(\mathrm{Background\,Subject\,Match}+\mathrm{Physical\,Realism}+\mathrm{Visual\,Quality}))$.
-To demonstrate the consistency of our evaluation metrics, we examined the correlation between different VLMs.
-We use the evaluated results for Nano Banana and GPT-Image-1, and examine the correlation coefficient between GPT’s and Gemini’s evaluation score.
-To ensure the reliability of the correlation coefficient, we divided the evaluation results into ten subsets and computed the average correlation across them.
+Omnigen2 [80] 0.2195 0.2927 0.0427 0.0793 0.2083 0.2500 0.1000 0.0778
+Kontext [41] 0.0488 0.1220 0.0183 0.0122 0.2500 0.3750 0.0556 0.1222
+Qwen-Image-Edit-2509 [79] 0.2683 0.2927 0.0488 0.1159 0.1250 0.2917 0.1111 0.1556
+DreamOmni2 [83] 0.5854 0.6585 0.5854 0.6280 0.5833 0.6667 0.5778 0.6333
+Nano Banana [22] 0.6829 0.7073 0.6463 0.5488 0.5000 0.5417 0.5556 0.5488
+GPT-Image-1 [56] 0.6829 0.7805 0.7195 0.7439 0.6250 0.6250 0.6889 0.6333
+D Comparison with Prior Benchmarks
+Table 1 shows that existing benchmarks do not provide systematic evaluation across diverse multi-reference conditions, support only a limited number of references, and fail to adequately account for heterogeneity among reference images.
+To illustrate that existing benchmarks are almost solved by state-of-the-art models such as Nano Banana [22] and GPT-Image-1 [56], we evaluate the image generation capabilities of these models using ImgEdit [91], the latest benchmark for image editing, and DreamOmni2 [83], the state-of-the-art benchmark for multi-reference image generation.
+The quantitative results show that recent closed-source, state-of-the-art models achieve consistently high scores, indicating that these benchmarks are nearing saturation and may soon be unable to meaningfully distinguish among high-performing models (Table 8 and Table 8).
+The evaluation results are taken from the official GitHub repository of ImgEdit [^1] https://github.com/PKU-YuanGroup/ImgEdit?tab=readme-ov-file and the results reported in the DreamOmni2 paper.
+Qualitative results suggest that, even in the additional “Hard” category of ImgEdit, most tasks are already solvable by advanced models such as Nano Banana (Figure 11).
+A similar tendency is observed for DreamOmni2 benchmark as well (Figure 11).
+Taken together, these findings indicate that current benchmarks for image editing and multi-reference image generation are approaching their limits in evaluating cutting-edge models.
+There is a clear need for more challenging benchmarks that better capture the capabilities and failure modes of the latest generation of image models.
+E Reliability and Cost of VLM Judges
+As described in Section 3.4, we employ VLM judges as a time- and cost-efficient proxy for human evaluation.
+In Section 4.4, we demonstrate that the VLM judge correlates strongly with human ratings, supporting our usage of VLM.
+Here, we provide a more detailed analysis of their reliability and the associated cost considerations.
+E.1 Cross-VLM Correlation
+To demonstrate the consistency of our evaluation metrics, we examined correlations among different VLMs.
+We use the evaluated results for Nano Banana and GPT-Image-1 and examine the correlation between GPT’s and Gemini’s evaluation scores.
+To assess the reliability of the correlation coefficient, we partitioned the evaluation results into 10 subsets and computed the mean correlation across them.
 Table 9 shows that the evaluated scores exhibit positive correlations, which demonstrates that our evaluation criteria are well-defined and consistent.
-Table 6: Correlation coefficients between GPT’s and Gemini’s evaluated scores. $\pm$ shows 95% confidence interval.
+Table 9: Correlation between GPT’s and Gemini’s evaluated scores. $\pm$ shows 95% confidence interval.
 Evaluation criteria Correlation coefficients
@@
 Background Subject Match 0.601 ± $\pm\,0.020$
-Physical Reallism 0.620 ± $\pm\,0.022$
+Physical Realism 0.620 ± $\pm\,0.022$
 Visual Quality 0.577 ± $\pm\,0.018$
 Overall 0.650 ± $\pm\,0.014$
+E.2 Self-Consistency Analysis
+We measure the standard error of the VLM judge scores across three different random seeds in Table 10.
+On the 1–10 rating scale, the observed standard errors are below 0.1, and we do not observe an increasing trend in standard error as the number of references increases, indicating that the VLM judge provides self-consistent evaluations for multi-reference image generation.
+Table 10: Standard error of total scores evaluated by the average of GPT and Gemini, for GPT-Image-1 and Nano Banana.
+Model 3-ref 4-ref 5-ref 6-ref 7-ref 8-ref
+GPT-Image-1 .0234.0234 .0138.0138 .0139.0139 .0153.0153 .0343.0343 .0244.0244
+Nano Banana .0260.0260 .0423.0423 .0033.0033 .0089.0089 .0062.0062 .0254.0254
+E.3 Sensitivity Analysis
 To further ensure the reliability of our evaluation metrics, we examined whether they can detect performance gains from fine-tuning on external image-editing datasets, thereby demonstrating that they align well with those used in other image-editing benchmarks.
@@
 We conducted this comparison using single-reference tasks because Qwen-Image-Edit supports only a single input image.
+E.4 Cost of VLM Judge
+Regarding evaluation costs, a single model evaluation costs approximately \$44 with GPT-5 and \$29 with Gemini 2.5 in January 2026.
+As an API-free alternative, we confirmed that Qwen3-VL [2] can be reliable enough (see Table 4).
+F Detailed Results
+F.1 Per-Task Evaluation
+Figure 15, Figure 16, and Figure 17 show the scores of each model across five evaluation metrics for each task, evaluated by GPT, Gemini, and their average, respectively.
+Taking the weighted average across five metrics, Figure 18 shows the total score of each model for single and two-reference tasks, evaluated by GPT, Gemini, and their average, respectively.
+For multi-reference tasks, Table 12, Table 13, and Table 14 show the total score, evaluated by GPT, Gemini, and their average, respectively.
+Overall, we observed a substantial gap in Instruction Alignment and Reference Consistency between the closed-source models and the other open-source models.
+In particular, GPT-Image-1 achieves significantly superior performance in Instruction Alignment, especially in style and text modification tasks in the two-reference task.
+Meanwhile, in Reference Consistency, Nano Banana achieves the highest scores in the single-reference task.
+Nano Banana also performs on par with GPT-Image-1 in Reference Consistency across background modification, subject replacement, subject addition, and makeup tasks, indicating its strong ability to leverage reference images, particularly when the number of references is small.
+On the other hand, Nano Banana performs worse than the other models in background modification tasks in Background Subject Match, Physical Realism, and Visual Quality.
+This result suggests that Nano Banana struggles with background modification.
 Table 11: Comparison of Qwen-Image-Edit and Qwen-Image-Edit-2509 on single-reference task.
@@
 single 6.332 7.499
-E Detailed Results
-E.1 Further Discussion of Per-Task Evaluation
-Figure 12, Figure 13, and Figure 14 show the scores of each model across five evaluation metrics for each task, evaluated by GPT, Gemini, and their average, respectively.
-Taking the weighted average across five metrics, Figure 15 shows the overall score of each model for single and two-reference tasks, evaluated by GPT, Gemini, and their average, respectively.
-For multi-reference tasks, Table 8, Table 9, and Table 10 show the overall score, evaluated by GPT, Gemini, and their average, respectively.
-Overall, we observed a substantial gap in Instruction Alignment and Reference Consistency between the closed-source models and the other open-source models.
-In particular, GPT-Image-1 achieves significantly superior performance in Instruction Alignment, especially in style and text modification tasks in the two-reference task.
-Meanwhile, in Reference Consistency, Nano Banana achieves the highest scores in the single-reference task.
-Nano Banana also performs on par with GPT-Image-1 in Reference Consistency, across background modification, subject replacement, subject addition, and makeup tasks, which indicates its strong ability to leverage reference images, particularly when the number of references is small.
-Additionally, Qwen-Image-Edit-2509, the most advanced open-source model, performs worse than the other models in background modification tasks in Background Subject Match and Visual Quality.
-This result suggests that it struggles with background modification.
-Figure 12: Scores of each model across evaluation metrics for each task by GPT.
-The horizontal axis denotes the scores for the five evaluation criteria, and the vertical axis denotes the number of reference images.
-Figure 13: Scores of each model across evaluation metrics for each task by Gemini.
-The horizontal axis denotes the scores for the five evaluation criteria, and the vertical axis denotes the number of reference images.
-Figure 14: Average scores of each model across evaluation metrics for each task by GPT and Gemini.
-The horizontal axis denotes the scores for the five evaluation criteria, and the vertical axis denotes the number of reference images.
-Figure 15: Overall scores of each model for single and two-reference tasks by GPT, Gemini, and their average.
-Table 8: Detailed per-task overall scores for the multi-reference tasks, evaluated by GPT.
+Table 12: Detailed per-task total scores for the multi-reference tasks, evaluated by GPT.
 The local, global, back, and object columns under X correspond to X–1 Objects + Local, X–1 Objects + Global, X–1 Objects + Background, and X Object, respectively.
@@
 GPT-image-1 6.16 6.08 5.68 5.90 5.38 6.09 5.59 5.94 4.96 5.82 5.03 5.15
-Table 9: Detailed per-task overall scores for the multi-reference tasks, evaluated by Gemini.
+Table 13: Detailed per-task total scores for the multi-reference tasks, evaluated by Gemini.
 The local, global, back, and object columns under X correspond to X-1 Objects + Local, X–1 Objects + Global, X–1 Objects + Background, and X Object, respectively.
@@
 GPT-image-1 3.66 4.66 3.46 4.02 4.18 4.58 3.24 3.52 3.28 4.07 2.91 2.96
-Table 10: Detailed per-task overall scores for each multi-reference task, evaluated by the average of GPT and Gemini.
+Table 14: Detailed per-task total scores for each multi-reference task, evaluated by the average of GPT and Gemini.
 The local, global, back, and object columns under X correspond to X–1 Objects + Local, X–1 Objects + Global, X–1 Objects + Background, and X Object, respectively.
@@
 GPT-image-1 4.91 5.37 4.57 4.96 4.78 5.33 4.42 4.73 4.12 4.94 3.97 4.05
-Figure 16: Overall scores for each difficult reference combination across models.
-Darker colors represent the average score of tasks that include the corresponding combination, while lighter colors indicate the average score of tasks that do not include it.
-Figure 17: Qualitative example for 4-Object Tasks.
-Figure 18: Qualitative example for 3-Object + Background Tasks.
-Figure 19: Qualitative example for 3-Object + Local Tasks.
-Figure 20: Qualitative example for 3-Object + Global Tasks.
-Figure 21: Qualitative example for Tasks with 8 references.
-E.2 Further Discussion of the Effect of the Number of References
-According to the per-metric scores in Figure 14 and the overall scores in Table 10, all models exhibit decreasing performance as the number of reference images increases.
+F.2 The Effect of the Number of References
+According to the per-metric scores in Figure 17 and the total scores in Table 14, all models exhibit decreasing performance as the number of reference images increases.
 For Instruction Alignment and Reference Consistency, GPT-Image-1 and Nano Banana achieve relatively high scores, whereas the open-source model, DreamOmni2, tends to approach the minimum score of 1 as the number of references increases.
@@
 As discussed in Section 4, this is because closed-source models prioritize adherence to the references and editing instructions over visual quality, while open-source models show a tendency to preserve visual quality but often ignore the references.
-In the finer-grained task-level evaluation, we found that even as the number of reference images increased, the X–1 Object + Global task did not exhibit score degradation in Background Subject Match, Physical Realism, or Visual Quality.
-This is because the task requires modifying the image to a specified style; therefore, even if the reference images come from different domains, which may degrade the visual quality, the final output converges to a unified style.
-In the results of Qwen-Image-Edit-2509, the model is trained to handle up to four reference images; therefore, when provided with five or more references, the input becomes out-of-distribution.
-This causes the model to produce perceptually invalid outputs, and then the scores approach the minimum value of 1.
-E.3 Further Discussion of Difficult Reference Combination
+In the finer-grained task-level evaluation, we found that, even as the number of reference images increased, the X–1 Object + Global task did not show severe decreases in Background Subject Match, Physical Realism, or Visual Quality scores.
+The task requires modifying the image to a specified style; therefore, even if the reference images come from different domains, which may degrade the visual quality, the final output often converges to a unified style.
+In the results of Qwen-Image-Edit-2509, the model may be trained to handle up to three reference images; therefore, when provided with four or more references, the input becomes out-of-distribution.
+This causes the model to produce perceptually invalid outputs, and then the scores approach the minimum value of 1 (see Figure 29).
+F.3 Difficult Reference Combination
 Figure 19 shows the scores of each model in difficult reference combination tasks across evaluation metrics.
-Cross-domain diversity   In cross-domain cases, samples with reference images from different domains (darker color) were evaluated as lower Reference Consistency scores, because all models aim to unify the domains during generation.
+Cross-domain diversity   In cross-domain cases, samples with reference images from different domains (darker color) received lower Reference Consistency scores because all models tend to distort individual reference attributes to achieve a consistent style.
 Scale and viewpoint differences   In the different scale and viewpoint cases, the Reference Consistency scores declined consistently across all models.
-According to the reasoning traces of the VLM, the primary reasons for deductions are the loss of fine details and changes in pose.
-This suggests that when models attempt to reproduce objects under different scales or viewpoints, they may fail to preserve fine details or may adjust the pose to achieve a more natural appearance.
-Rare concept references   In the rare concept case, because the reference images often contain subjects at a large scale, instances were observed in which the model reproduced this scale without adjustment, even when it was too large.
-For this reason, rare concept samples tend to receive lower scores on the Physical Realism.
+This suggests that when models attempt to reproduce objects at different scales or from different viewpoints, they may fail to preserve fine details or adjust the pose to achieve a more natural appearance.
+Rare concept references   In the rare-concept case, the model may struggle to handle uncommon subjects relative to more familiar ones.
+Because the reference images often depict these subjects at a large scale, the model tends to reproduce this scale without appropriate adjustment.
+This suggests that, unlike common concepts, the model may find it more difficult to flexibly control attributes such as size when dealing with rare concepts.
+As a result, rare concept samples tend to receive lower Physical Realism scores.
 Multilingual references   In the multilingual text rendering case, we observed a consistent decline across all metrics except Reference Consistency.
-This primarily occurs because models have limited text-rendering capability in languages other than English, leading to failures in cross-language conversion.
-As a result, instruction alignment and visual quality become lower.
-However, scores dropping in GPT-Image-1 were lower than those of others, since it exhibits comparatively stronger text-rendering ability in non-English languages.
-E.4 Agentic Inference
+This may occur because models have limited text-rendering capability in languages other than English, leading to failures in cross-language conversion.
+As a result, instruction alignment and visual quality decrease.
+However, GPT-Image-1 exhibits a smaller decline than other models, owing to its comparatively stronger text-rendering ability in non-English languages.
+F.4 Risks of Model Bias in Synthetic References
+Synthetic references generated by closed models (i.e., Nano Banana, GPT-Image-1) may introduce in-distribution advantages when those same models are also evaluated on the benchmark.
+To verify this, we divided the benchmark into three subsets based on the source of the reference images: those from GPT-Image-1, those from Nano Banana, and those from both.
+The mean scores across all subsets remain within their respective 99% confidence intervals, indicating no statistically significant bias (Figure 12).
+Figure 12: Analysis of potential in-distribution bias. Error bars indicate the 99% confidence intervals. For image generation models, we use the versions available as of January 2026.
+F.5 Potential Conflict in Cross-Domain Task
+One might expect that preserving subject details (reference consistency) and maintaining a coherent scene (background-subject match) are inherently at odds in cross-domain tasks—e.g., placing a photorealistic person into an anime background.
+However, our evaluation shows this is not always the case.
+As shown in Figure 13 (left), a generated image can satisfy both criteria, achieving a GPT-5 score of 7 for reference consistency and 9 for background-subject match.
+To further verify this, we created cross-domain editing tasks with explicit instructions that prioritize one criterion over the other for 20 randomly selected prompts.
+Table 15 shows that Nano Banana can improve the prioritized criterion while maintaining the other criterion.
+Table 15: Comparison of the mean scores when prioritizing either reference consistency or background–subject match for the subset generated by Nano Banana, latest version as of January 2026.
+Prompt
+Reference
+Consistency
+Background–
+Subject Match
+Original 3.58 3.67
+Consistency priority 4.17 3.25
+Matching priority 3.40 4.03
+Figure 13: Left: A sample generated by GPT-Image-1 on a cross-domain task. Right: Task example of our benchmark directly relevant to real-world use cases, such as advertising.
+F.6 Alignment with Real-World Use Cases
+As shown in Figure 13 (right), MultiBanana is directly relevant to real-world use cases, such as advertising.
+Our benchmark development using synthetic data also aligns with the community standard.
+For example, ImgEdit [91] similarly employs GPT-4o for instruction generation, which also aims to develop “practical and powerful tools for real-world applications.”
+F.7 Agentic Inference
 Figure 14: Example of changes in the Iterative Prompt Refinement (IPR) framework.
 Comparison and evaluation results between Nano Banana and GPT on the task of generating images from 8 object references according to the instruction prompt.
-We introduced three agentic frameworks, Iterative Prompt Refinement (IPR), Context-Aware Feedback Generation (CAFG), and Selective Reference Adaptation (SRA).
+We introduced three agentic frameworks: Iterative Prompt Refinement (IPR), Context-Aware Feedback Generation (CAFG), and Selective Reference Adaptation (SRA).
 Here, we refer to $\mathrm{Gen}$ as the Generator, which produces an image, and $\mathrm{Plan}$ as the Planner, which updates the instruction prompt (and selects reference images) for the next step.
@@
 For GPT, we use GPT-Image-1 as the Generator and GPT-5 as the Planner.
-For Gemini, we use Nano Banana as the Generator and Gemini-2.5-Flash as the Planner.
-E.4.1 Iterative Prompt Refinement (IPR)
-IPR framework is formulated as follows. The Planner refines the prompt based on the generation result from the previous step.
+For Gemini, we use Nano Banana as the Generator and Gemini 2.5 Flash as the Planner.
+F.7.1 Iterative Prompt Refinement (IPR)
+The IPR framework is formulated as follows. The Planner refines the prompt based on the generation result from the previous step.
 $G^{t+1}=\mathrm{Gen}(P^{t},\{R_{i}\}_{i\in[I]},\varnothing),$
 $P^{t+1}=\mathrm{Plan}(P^{t},\{R_{i}\}_{i\in[I]},G^{t+1}).$
-E.4.2 Context-Aware Feedback Generation (CAFG)
-CAFG framework is formulated as follows. The Generator generates the image based on the generation result (context) from the previous step.
+F.7.2 Context-Aware Feedback Generation (CAFG)
+The CAFG framework is formulated as follows. The Generator generates the image based on the generation result (context) from the previous step.
 $G^{t+1}=\mathrm{Gen}(P^{t},\{R_{i}\}_{i\in[I]},\bm{G^{t}}),$
 $P^{t+1}=\mathrm{Plan}(P^{t},\{R_{i}\}_{i\in[I]},G^{t+1}).$
-E.4.3 Selective Reference Adaptation (SRA)
-SRA framework is formulated as follows.
+F.7.3 Selective Reference Adaptation (SRA)
+The SRA framework is formulated as follows.
 The Planner selects only the reference images that should be improved based on the generation result from the previous step, and the Generator receives these as context.
-SRA is expected to reduce task complexity for the agent by adaptively decreasing the number of references.
+SRA is expected to reduce agents’ task complexity by adaptively decreasing the number of references.
 $G^{t+1}=\mathrm{Gen}(P^{t},\{R_{i}\}_{\bm{i\in U^{t}}},G^{t}),$
@@
 where $U^{t}$ is the index set of reference images that are insufficiently reflected in the generated image $G^{t}$.
-E.4.4 Experiments
-We set the maximum number of steps $t$ as $3$ and we evaluate three agentic frameworks—IPR, CAFG, and SRA—using Gemini, and the IPR framework using GPT.
+F.7.4 Experiments
+We set the maximum number of steps $t$ to $3$ and evaluate three agentic frameworks—IPR, CAFG, and SRA—using Gemini, and the IPR framework using GPT.
 Figure 14 illustrates an example of step-wise improvement in the IPR framework.
-Figure 23, Figure 24, Figure 25, and Figure 26 present the evaluation results of multi-reference generations at each step, averaged across judgments from Gemini-2.5-Flash and GPT-5, broken down by category.
+Figure 20, Figure 21, Figure 22, and Figure 23 present the evaluation results of multi-reference generations at each step, averaged across judgments from Gemini 2.5 Flash and GPT-5, broken down by category.
 Figure 24 presents the results for single and two-reference generations.
-Nano Banan (Gemini) shows modest improvements in Physical Realism and Visual Quality across refinement steps, while Instruction Alignment and Reference Consistency either remain unchanged or deteriorate.
+Nano Banana (Gemini) shows modest improvements in Physical Realism and Visual Quality across refinement steps, while Instruction Alignment and Reference Consistency either remain unchanged or deteriorate.
 In contrast, GPT demonstrates consistent improvements across all categories.
 This suggests that Gemini’s planner progressively loses information from the original prompt as refinement steps proceed.
-IPR Framework with Gemini
-Figure 23: Detailed results of multi-reference image generation using the IPR framework with Gemini.
-CAFG Framework with Gemini
-Figure 24: Detailed results of multi-reference image generation using the CAFG framework with Gemini.
-SRA Framework with Gemini
-Figure 25: Detailed results of multi-reference image generation using the SRA framework with Gemini.
-IPR Framework with GPT
-Figure 26: Detailed results of multi-reference image generation using the IPR framework with GPT.
-Figure 27: Detailed results of single and two-reference image generation.
-F Extended Related Works
-F.1 Controllable Text-to-Image Generation
-Diffusion models have achieved remarkable progress in generative modeling, demonstrating exceptional performance, particularly in high-fidelity image synthesis.
-Representative systems, such as Stable Diffusion [rombach2022ldm, podell2024sdxl, esser2024sd3], FLUX [flux2024, labs2025flux1kontext], DALL-E [ramesh2022dalle], and Imagen [saharia2022imagen, imagenteamgoogle2024imagen3], have demonstrated strong text-to-image generation capabilities, establishing a controllable and scalable foundation for image generation.
-To enhance controllability, models such as ControlNet [zhang2023controlnet] and T2I-Adapter [mou2023t2i-adapter] introduced external conditioning modules, enabling image-conditioned generation.
-Recent large pretrained diffusion models with textual conditioning, such as Stable Diffusion [rombach2022ldm, podell2024sdxl, esser2024sd3] and FLUX [flux2024, labs2025flux1kontext], enable the synthesis of novel scene images while preserving the identity depicted in the reference images.
-Building on these developments, approaches based on fine-tuning (e.g., DreamBooth [ruiz2022dreambooth]) and adapter-based methods (e.g., IP-Adapter [ye2023ip-adapter]) have been actively explored to address this challenge [jiang2025infiniteyou, gal2023an, brooks2023instructpix2pix, mokady2023null, tan2025ominicontrol, NEURIPS2023_602e1a5d].
-Additionally, training-free approaches (e.g., pix2pix-zero [parmar2023zero] and prompt-to-prompt [hertz2023prompttoprompt]) have also been widely proposed [Cao_2023_ICCV, Huberman-Spiegelglas_2024_CVPR, Wallace_2023_CVPR, Tumanyan_2023_CVPR, kulikov2025flowedit, miyake2025negative-prompt-inversion].
+G Implementations
+API endpoints: For VLM evaluation, we used stable API versions: gpt-5-2025-08-07 and gemini-2.5-flash.
+For image generation models, we use the latest available API versions as of November 2025, unless otherwise noted.
+H Extended Related Works
+H.1 Controllable Text-to-Image Generation
+Stable Diffusion [65, 60, 15], FLUX [42, 41], DALL-E [63], and Imagen [67, 32], have demonstrated strong text-to-image generation capabilities, establishing a scalable foundation for the task.
+To enhance controllability, models such as ControlNet [94] and T2I-Adapter [52] introduced external conditioning modules, enabling image-conditioned generation.
+Additionally, training-free approaches (e.g., pix2pix-zero [59] and prompt-to-prompt [25]) have also been widely proposed [6, 31, 75, 74, 39, 49].
 These advancements collectively demonstrate that diffusion models are becoming increasingly flexible and adaptable across diverse conditional generation settings.
-In recent years, research on unified image generation has gained momentum, aiming to handle multiple image generation tasks within a single model.
-For example, ChatGPT-Image-1 [openai2025gpt4oimage] and Nano Banana [google2025nanobanana] are capable of jointly processing text and image inputs, enabling integrated image generation and editing within a unified framework.
-Similarly, the open-source models, such as Qwen-Image [wu2025qwenimagetechnicalreport] and FLUX.1-Kontext-dev [labs2025flux1kontext], built upon a robust multimodal large language model backbone, demonstrate high-quality and flexible image generation and editing performance.
-Beyond these, numerous open-source unified generative models continue to emerge [xiao2024omnigen, xia2025dreamomni, wu2025omnigen2, xia2025dreamomni2].
-With the advent of multimodal foundation models capable of jointly handling text and images [google2025nanobanana, openai2025gpt4oimage], a new task has recently become feasible: multi-reference-based image generation.
-In this task, the model receives multiple reference images along with textual instructions and generates a new scene while preserving the identity of the reference images.
-However, despite its novelty and potential, a reliable benchmark for evaluating this innovative task has yet to be established, leaving it as an essential open challenge for future research.
-F.2 Benchmarks for Reference-Based Generation
+H.2 Benchmarks for Reference-Based Generation
+Reference-based image generation is closely related to industrial applications such as content production [66, 90, 78, 96, 88, 37], advertising [33, 28, 51], and fashion design [98, 11, 16, 16, 12, 86, 36].
 Among reference-driven image generation tasks, the most widely recognized one is image editing.
-Image editing refers to the task of generating an edited image based on a single reference image and a textual instruction describing how the edit should be performed.
-Representative benchmarks such as MagicBrush [zhang2023magicbrush] and EMU-Edit [sheynin2024emuedit] have extended task-specific evaluations but still rely heavily on similarity-based metrics.
-SmartEdit [huang2024smartedit] supports editing of complex scenes, but does not sufficiently cover more general settings.
-Similarly, I2E-Bench [ma2024i2ebench] employs GPT-4o to provide human-aligned evaluations across a variety of editing tasks, yet it uses different metrics for each task and therefore fails to capture the shared characteristics of editing as a whole.
-More recently, ImgEdit [ye2025imgedit] has enabled the evaluation of multi-turn editing tasks, where image generation and editing are alternated interactively.
-On the other hand, one of the earliest benchmarks for reference-based image generation beyond editing is the DreamBooth [ruiz2022dreambooth] dataset.
-However, such benchmarks are still limited to generation tasks conditioned on a single reference image and thus cannot handle more complex or compositional settings.
-In recent years, following advances demonstrated by GPT-Image-1 [openai2025gpt4oimage] and Nano Banana [google2025nanobanana], increasing attention has been directed toward multi-reference image generation, where multiple reference images are used jointly for generation.
-Benchmarks such as OmniContext [wu2025omnigen2] and DreamOmni2 [xia2025dreamomni2] have been proposed to evaluate this task.
-Nevertheless, these benchmarks encompass only a small number of images and task types, thereby limiting the scope of the evaluation.
-Moreover, despite being designed for multi-reference generation, they do not evaluate qualitative differences or relationships between reference images, and their setups remain similar to those of conventional single-image editing benchmarks.
-G Prompts
-G.1 Prompts of Agentic Framework
+RealEdit [71] focuses on single-image editing and provides an empirical analysis of real-world image-editing use cases.
+SpotEdit [20] proposes a benchmark for visually guided image editing.
+These works focus on benchmarking single-reference image edition/generation, while we focus on multi-reference image generation.
+In recent years, increasing attention has been directed toward multi-reference image generation, in which multiple reference images are jointly used to generate a single image.
+MultiRef-Bench [9] studies controllable image generation with multiple visual references, but our benchmark outperforms it by providing 3,769 samples (compared to 1,990), supporting up to 8 references (compared to 6), and using raw images (compared to bounding boxes or masks), making it more practical.
+H.3 Instruction-Following in LLMs
+The task of generating an image that faithfully reflects multiple references is analogous to instruction following in LLMs [97, 95, 46, 14, 76], in which models must simultaneously satisfy multiple constraints.
+In the text domain, recent studies have revealed that LLMs struggle to follow all given instructions as the number of constraints grows, often ignoring or inadequately addressing a subset of them [35, 24, 23, 40].
+Multi-reference image generation poses a parallel challenge: as the number of reference images increases, models must reconcile potentially conflicting visual cues while remaining faithful to every reference.
+I Limitations and Discussions
+Our benchmark focuses on static image generation and does not address video generation or editing [27, 21, 55], where temporal consistency across frames introduces additional challenges for reference-based conditioning [29, 10, 1].
+Moreover, our prompts describe object placement in natural language (e.g., ”on the left,” ”in the foreground”), which is inherently ambiguous.
+Combining reference images with structured layout inputs [94, 89, 45, 61] could enable finer-grained control over subject composition.
+J Prompts
+J.1 Prompts of Agentic Framework
 Prompt of the Generator in the IPR framework {prompt}{reference_image_files}
@@
 Emphasize the color palette and texture details.” } {reference_image_files}{previously_generated_image_file}
-G.2 Prompts of VLM as Judge
+J.2 Prompts of VLM as Judge
 […]
@@
 […]
+Figure 15: Scores of each model across evaluation metrics for each task by GPT.
+The horizontal axis denotes the scores for the five evaluation criteria, and the vertical axis denotes the number of reference images.
+Figure 16: Scores of each model across evaluation metrics for each task by Gemini.
+The horizontal axis denotes the scores for the five evaluation criteria, and the vertical axis denotes the number of reference images.
+Figure 17: Average scores of each model across evaluation metrics for each task by the average of GPT and Gemini.
+The horizontal axis denotes the scores for the five evaluation criteria, and the vertical axis denotes the number of reference images.
+Figure 18: Total scores of each model for single and two-reference tasks by GPT, Gemini, and their average.
+Figure 19: Total scores for each difficult reference combination across models.
+Darker colors represent the average score for tasks that include the corresponding combination, whereas lighter colors indicate the average score for tasks that do not include it.
+IPR Framework with Gemini
+Figure 20: Detailed results of multi-reference image generation using the IPR framework with Gemini.
+CAFG Framework with Gemini
+Figure 21: Detailed results of multi-reference image generation using the CAFG framework with Gemini.
+SRA Framework with Gemini
+Figure 22: Detailed results of multi-reference image generation using the SRA framework with Gemini.
+IPR Framework with GPT
+Figure 23: Detailed results of multi-reference image generation using the IPR framework with GPT.
+Figure 24: Detailed results of single and two-reference image generation using the agentic frameworks.
+Figure 25: Qualitative example for 4-Object Tasks.
+Figure 26: Qualitative example for 3-Object + Background Tasks.
+Figure 27: Qualitative example for 3-Object + Local Tasks.
+Figure 28: Qualitative example for 3-Object + Global Tasks.
+Figure 29: Qualitative example for Tasks with 8 references.